### v1

# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

Abstract
We describe a novel subgradient following apparatus for calculating the optimum of convex problems with variational penalties. In this setting, we receive a sequence $y_{i},\ldots,y_{n}$ and seek a smooth sequence $x_{1},\ldots,x_{n}$. The smooth sequence attains the minimum Bregman divergence to an input sequence with additive variational penalties in the general form of $\sum_{i}g_{i}(x_{i+1}-x_{i})$. We derive, as special cases of our apparatus, known algorithms for the fused lasso and isotonic regression. Our approach also facilitates new variational penalties such as non-smooth barrier functions. We next derive and analyze multivariate problems in which ${\mathbf{x}}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ and variational penalties that depend on $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. The norms we consider are $\ell_{2}$ and $\ell_{\infty}$ which promote group sparsity. Last but not least, we derive a lattice-based subgradient following for variational penalties characterized through the output of arbitrary convolutional filters. This paradigm yields efficient solvers for problems in which sparse high-order discrete derivatives such as acceleration and jerk are desirable.
1 Sub-Introduction
Rather than descending into a wordy introduction conceived by ESL writers which was then passed through a spell checker and further beautified by an extra-fine-tuned language model, let us delve right into a simple gateway for this paper by examining the so called fused lasso [16]. This problem is concerned with the following optimization task,
$\min_{{\mathbf{x}}\in\mathbb{R}^{n}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n}\lambda_{i}|x_{i+1}-x_{i}|~{}~{}.$ (1)
The first part of the paper is concerned with an abstract unified treatment that generalizes the setting of the fused lasso to optimization problems of the form,
$\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}g_{i}(x_{i}-x_{i+1})~{}~{}.$
Here, both $h_{i}:\mathbb{R}\to\mathbb{R}_{+}$ and $g_{i}:\mathbb{R}\to\mathbb{R}_{+}$ are convex in their single argument. In addition to retrieving, as special cases, both the fused-lasso and isotonic regression, our algorithm apparatus can be used with non-smooth variational penalties for $g_{i}$. To illustrate its power we derive an entirely new variant with barrier functions serving as the skeleton for variational penalty. We fondly refer to this refreshingly new variant as the Untangled Lariat (if Rob could so can we).
We explored uncharted water in which each element of the input sequence is a vector and thus need to find an sequence ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ where ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ for all $i\in[n]$. In order to obtain variational sparsity in the multivariate case we replace the absolute value $|x_{i+1}-x_{i}|$ with norm penalties over the differences $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. As a starting point we derive an aesthetically succinct closed form for the squared $2$-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$. We use the closed form solution to build surrogate functions for the $2$-norm itself, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$, and the infinity-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$ yielding an efficient iterative solver for the problem,
$\min_{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}}\>\sum_{i=1}^{n}h_{i}({\mathbf{x}}_{i})+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i+1}\|_{p}~{}\mbox{ where }~{}p\in\{2,\infty\}.$
For brevity of the derivation we focus on the case where $h_{i}({\mathbf{x}}_{i})=\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}$. To the best of our minute yet tangible knowledge these settings and their analyses are new and provide new data-science and machine learning tools for multivariate temporal problems with variational sparsity.
In the last part we incorporate a common construction from signal processing and penalize the output of the convolution filter applied to the input sequence $y_{1},\ldots,y_{n}$. We provide efficient apparatus for finding the optimum of,
$\min_{\mathbf{x}}\sum_{i}h_{i}(x_{i})+\sum_{i}\big{|}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{|}~{},$
where $\boldsymbol{\tau}\in\mathbb{R}^{k}$ is an arbitrary filter of $k-1$ free parameters. We introduce a subgradient following mechanism over a $k$-dimensional lattice requires $\mathcal{O}(n^{k})$ to construct in the worst case. In applications where $\mathbf{y}$ represents a temporal sequence, the variational penalty $x_{i+1}-x_{i}$ can be viewed as promoting smoothed output with regions of zero discrete velocity. Our general framework can be used as a smoothing apparatus for specific problems in which we seek regions of zero acceleration, zero jerk, and other forms of finite differences.
Before diving in, we do want to pay some homage to previous work and make connections to our framework. The core of related research preceding this paper can be traced back to the work on the fused lasso [16]. The setting of the fused lasso has been revisited numerous times and found various venues for applications, see for instance [7, 8], though any attempt to be exhaustive here is going to injustice to the plethora of applications that employ the fused lasso. Several authors [4, 11, 12] described and analyzed a linear time algorithm for solving the fused lasso problem. In addition, alternative algorithms such as the ones described in [1, 3] do not guarantee linear time yet demonstrate excellent empirical results. The most relevant paper to our work is [12] where the authors follow the derivation in [11] to generalize the algorithm to non-smooth unary functions and to trees. Their procedure is an instance of message passing algorithm [6]. The paper [12] also considers different unary functions subject to a pairwise term. The term includes the fused lasso penalty as a special case. The lens of the work is by nature probabilistic and does not discuss general pairwise penalties. We encapsulate and generalize the algorithmic work above by providing a new subgradient following algorithm for general convex unary functions and pairwise terms.
Our approach is morally different than the above work by employing the Fenchel dual of the primal problem, a view that was advocated by Shalev-Shwartz [15] in the context of online learning. This view also facilitates a simple yet nonetheless effective generalization to Bregman divergences [2] as the objective at no additional computation cost and, as mentioned above, is equally applicable to non-smooth variational penalties. We conclude with a short discussion on extensions and generalizations of this work such as incorporating the framework from [12] for general acyclic graphs to our setting. While left for future research, we believe that the approximation for general graphs of [13], that applies 1D fused lasso solver over induced chains can be applied as well. We provide in an appendix Python code that implements subgradient following procedures for many of the settings discussed in this paper.
2 Abstract Subgradient Following Algorithm
Let us now consider a general minimization with sequential penalties of the form,
$U=\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}g_{i}(x_{i}-x_{i+1})~{}~{}.$ (2)
Throughout the $h_{i}$ and $g_{i}$ are convex and lower semi-continuous. To remind the reader, the Fenchel dual of a function $f$ is defined as
$f^{*}(\alpha)=\sup_{x\in\mathrm{dom}(f)}\big{\{}\alpha x-f(x)\big{\}}~{}.$
The domain of the functions we consider is closed, compact, and convex and thus we can replace the supremum with a maximum. We next apply Fenchel-Young inequality to each $g_{i}$,
$g_{i}(x_{i}-x_{i-1})+g_{i}^{*}(\alpha)\geq\alpha(x_{i}-x_{i-1})~{},$
to bound $U$ from below,
$U$$\geq$ $\min_{{\mathbf{x}}}\,\max_{\boldsymbol{\alpha}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]$
$=$ $\min_{x_{1}}\,\max_{\alpha_{1}}\,\min_{x_{2}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]$
$=$ $\max_{\boldsymbol{\alpha}}\,\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]~{},$
where we used successive application of the minimax theorem in the two equalities above. We confine ourselves to settings for which strong duality holds thus,
$U$ $=$ $\min_{x_{1}}\,\max_{\alpha_{1}}\,\min_{x_{2}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]$
To find the optimum we define a sequence of partial single variable functions defined as,
$f_{i}(x_{i})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(x_{j}-x_{j+1})-g_{j}^{*}(\alpha_{j})\right]~{}.$ (3)
For boundary conditions we let $\forall x:f_{n}(x)=0$. In addition, we introduce the auxiliary function $\tilde{f}_{i}(x_{i})=f_{i}(x_{i})+h_{i}(x_{i})$. Using these definitions we introduce a summation-free form for $f_{i}$,
$f_{i}(x_{i})$ $=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+1}(x_{i+1})+h_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$ (4)
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{\tilde{f}_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})-\tilde{f}_{i+1}^{*}(\alpha_{i})~{}.$
Note that to find $f_{i}(x_{i})$ we need to obtain the minimum of $\tilde{f}_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}$ with respect to $x_{i+1}$. Using this abstraction, obtaining the optimum for $x_{i+1}$ requires “only” the knowledge of $\alpha_{i}$. Analogously, to calculate $f_{i}$ itself we need to find the optimum of $\alpha{}x_{i}-g_{i}^{*}(\alpha)-\tilde{f}_{i+1}^{*}(\alpha)$ with respect to $\alpha$ which gives us $\alpha_{i}$. This in turn only requires the knowledge of $x_{i}$. To recap, the optimization sub-problem for $x_{i+1}$ solely requires the knowledge of $\alpha_{i}$ whereas the optimization sub-problem for $\alpha_{i}$ solely requires the knowledge of $x_{i}$. Let $\mathsf{x}_{1},\mathsf{a}_{1},\mathsf{x}_{2},\mathsf{a}_{2},\ldots,\mathsf{a}_{n-1},\mathsf{x}_{n}$ be a sequence of locally-admissible pairs $\mathsf{x}_{i},\mathsf{a}_{i}$ and $\mathsf{a}_{i},\mathsf{x}_{i+1}$ where the full sequence is not necessarily globally optimal. The subgradient conditions for pair-restricted optimality define two admissible sets,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1})\big{\}}$ (5)
$S_{i}$ $=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})\big{\}}~{}.$ (6)
Given $S_{i+1}=\left\{(\mathsf{x}_{i+1},\mathsf{a}_{i+1}):\mathsf{a}_{i+1}\in\partial f_{i+1}(\mathsf{x}_{i+1})\right\}$ we perform an expansion,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i+1}(\mathsf{x}_{i+1})+\partial h_{i+1}(\mathsf{x}_{i+1})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}-\mathsf{a}_{i+1}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i+1})\in S_{i+1}\big{\}}~{}.$
We next derive a recursive map $S_{i+1}\mapsto S_{i}$ by going through $S_{i+\frac{1}{2}}$. From the representation of $f_{i}(x_{i})$ as given by Eq. (16) we get that the pair $(\mathsf{x}_{i},\mathsf{a}_{i})$ is admissible when
$\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial\tilde{f}_{i+1}^{*}(\mathsf{a}_{i})~{}.$
Starting with the definition of $S_{i}$ and expanding it based on the above property we get,
$S_{i}$ $=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial{\tilde{f}_{i+1}}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}.$
For the last equality we used the fact that,
$(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}~{}\Leftrightarrow~{}\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial\tilde{f}^{*}_{i+1}(\mathsf{a}_{i})~{}.$
In summary, we obtain the following aesthetically succinct recursion,
$\begin{array}[]{lcl}S_{i+\frac{1}{2}}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{a}-\mathsf{a}^{\prime}\in\partial h_{i+1}(\mathsf{x})\;;\;(\mathsf{x},\mathsf{a}^{\prime})\in S_{i+1}\right\}\\ S_{i}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;(\mathsf{x}^{\prime},\mathsf{a})\in S_{i+\frac{1}{2}}\right\}~{}.\end{array}$
To kick-start the recursive procedure we set $S_{n}$ as a boundary condition,
$S_{n}\equiv\big{\{}(\mathsf{x},0):\mathsf{x}\in\text{dom}(h_{n})\big{\}}~{}.$
We end the recursion upon reaching $S_{\frac{1}{2}}$ and then set $x_{1}$ to be the left argument of the pair $(\mathsf{x},0)\in S_{\frac{1}{2}}$. Once $x_{1}$ is inferred we can read out the rest of the sequence, $x_{i}$ for $i>1$, by unravelling the recurrence,
[…]
We henceforth refer to the above general schema as well as the following derived algorithms as Nariat for a shorthand to uNtangled Lariats.
3 Derived Algorithms
In this section we apply the subgradient following framework to specific settings. We first show that the fused lasso is a simple instance. We also derive other variational (fused) penalties, including barrier penalties that were not analyzed before. We also show that the squared distance between ${\mathbf{x}}$ and $\mathbf{y}$ can be generalized at no additional computational cost to separable Bregman divergences.
3.1 Revisiting the Fused Lasso
To recover the Fused Lasso algorithm from the set following schema we choose, $h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}$ and $g_{i}(\delta)=\lambda_{i}|\delta|$. Using basic calculus these choices imply that $\nabla h_{i}(x)=x-y_{i}$ and $g_{i}^{*}(\alpha)=\mathbf{1}_{|\alpha|\leq\lambda_{i}}$. For the domain of $g_{i}^{\star}$, namely $|\alpha|\leq\lambda_{i}$, its subgradient is,
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=-\lambda_{i}\\ 0&|\alpha|<\lambda_{i}\\ [0,\infty)&\alpha=\lambda_{i}\end{cases}~{}.$
We next show that $\forall i\in[n]$, there exists a mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$, where $\mathsf{a}_{i}(\cdot)$ is a continuous non-decreasing piecewise linear function in $\mathsf{x}$, which serves as a succinct representation of $S_{i}=\{(\mathsf{x},\mathsf{a}_{i}(\mathsf{x}))\}$. The mapping takes the following recursive form,
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}.$ (7)
First, let us define $a_{n}(\mathsf{x})\equiv 0$ which adheres with the definition of $S_{n}$ from previous section. Next, we show that the form of $S_{i}$ holds inductively by using the definitions of $S_{i}$ and $S_{i+\frac{1}{2}}$ from previous section. Assume that Eq. 7 holds true for $S_{i+1}$. The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$S_{i+\frac{1}{2}}=\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):\mathsf{x}\in\mathsf{R}\}~{}.$
For convenience we define $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. In order to evaluate,
$S_{i}=\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})\}$
we need to analyze $\Delta\mathsf{x}=\mathsf{x}-\mathsf{x}^{\prime}$ for the admissible set of values of $\mathsf{a}$ under the map $g_{i}^{*}$ which is $[-\lambda_{i},\lambda_{i}]$. Since $\mathsf{a}_{i+1}(\cdot)$ is continuous and non-decreasing, we have $\mathsf{a}_{i}^{\prime}(.)$ is strictly-increasing and bijective. Therefore, there exists two unique boundary values $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=-\lambda_{i}$ and $z_{i}^{+}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{+})=+\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})$ such that $-\lambda_{i}<\mathsf{a}<\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=\{0\}$. Thus, $\mathsf{x}=\mathsf{x}^{\prime}$ and both are in $(z_{i}^{-},z_{i}^{+})$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\in(z_{i}^{-},z_{i}^{+})$ yields,
$S_{i}\cap\{\mathsf{x}:z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$ $=$ $\{(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x})):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$
$=$ $\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}~{}.$
We last need to address the cases $\mathsf{a}=\pm\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(z_{i}^{+})=\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=[0,\infty)$. Thus, for any $\mathsf{x}\geq z_{i}^{+}$ we have $\mathsf{x}-z_{i}^{+}\in\partial g^{*}_{i}(\lambda_{i})$ and therefore $(\mathsf{x},\lambda_{i})\in S_{i}$. Analogously, for $\mathsf{a}=-\lambda_{i}$ which implies that for $\mathsf{x}\leq z_{i}^{-}$ it holds that $\mathsf{x}-z_{i}^{-}\in\partial g^{*}_{i}(-\lambda_{i})\Rightarrow(\mathsf{x},-\lambda_{i})\in S_{i}$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\geq z_{i}^{+}$ and $\mathsf{x}\leq z_{i}^{-}$ yields,
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\geq z_{i}^{+}\}$ $=$ $\{(\mathsf{x},\lambda_{i}):\mathsf{x}\geq z_{i}^{+}\}$
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},-\lambda_{i}):\mathsf{x}\leq z_{i}^{-}\}~{}.$
Since $\mathsf{a}_{i}^{\prime}(\cdot)$ is increasing, we have for $\mathsf{x}\geq z_{i}^{+}\Rightarrow\mathsf{a}_{i}^{\prime}(\mathsf{x})\geq\lambda_{i}$ and $\mathsf{x}\leq z_{i}^{-}\Rightarrow\mathsf{a}_{i}^{\prime}(\mathsf{x})\leq-\lambda_{i}$. Therefore, we established that
$S_{i}=\left\{\Big{(}\mathsf{x},\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{-\lambda_{i}}^{+\lambda_{i}}\Big{)}:\mathsf{x}\in\mathsf{R}\right\}~{}.$
Figure 1: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for the fused lasso.
3.2 Sparse Fused Lasso
We next combine both the standard lasso and fusion penalties. Namely, we choose
$h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}+\beta_{i}|x|~{}.$
and $g_{i}(\delta)=\lambda_{i}|\delta|$. This choice implies that
$\partial h_{i}(x)=\begin{cases}x-y_{i}+\beta_{i}&x>0\\ [-y_{i}-\beta_{i},-y_{i}+\beta_{i}]&x=0\\ x-y_{i}-\beta_{i}&x<0\end{cases}~{}.$
Generalizing the fused lasso, there exists a succinct representation for the sparse fused lasso of the form
$S_{i}=\{(\mathsf{x},\mathsf{a}_{i}^{+}(\mathsf{x})):\mathsf{x}>0\}\,\cup\,\{(0,\mathsf{a}):\mathsf{a}_{i}^{-}(0)\leq\mathsf{a}\leq\mathsf{a}_{i}^{+}(0)\}\,\cup\,\{(\mathsf{x},\mathsf{a}_{i}^{-}(\mathsf{x})):\mathsf{x}<0\}~{}.$
The locally-admissible set consists of three disjoint subsets induced by the sign of the first argument $\mathsf{x}$. It is immediate to verify that the subgradient at any value but zero is unique. Hence, the mappings for positive and negative numbers are well-defined and take similar recursive forms to that of the fused lasso,
$\mathsf{a}_{i}^{\pm}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}^{\pm}(\mathsf{x})+\mathsf{x}-y_{i+1}\pm\beta_{i+1}\big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}.$
The sole value of that requires further attention is at $\mathsf{x}=0$. The admissible interval at zero is determined by the two boundary values, $\mathsf{a}_{i}^{-}(0)$ and $\mathsf{a}_{i}^{+}(0)$. To facilitate an efficient implementation, we store the value of $\Delta_{i}=\mathsf{a}_{i}^{+}(0)-\mathsf{a}_{i}^{-}(0)$. Before performing the rectification, we let $\Delta_{i}\leftarrow\Delta_{i-1}+2\beta_{i}$. Analogous to the fused lasso, we traverse the segments of $\mathsf{a}_{i}(\cdot)$ left-to-right until we reach the segment encapsulating $-\lambda_{i}$. The only difference is that we increase the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero. Analogously, we decrease the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero when traversing right-to-left. After obtaining the critical points $z_{i}^{\pm}$, we update $\Delta_{i}$ accordingly. In each iteration, we incur constant additional costs and thus the total runtime remains intact as $\mathcal{O}(n)$.
Figure 2: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for the sparse fused lasso.
3.3 Bregman Divergences
Before proceeding with less explored fused regularizations, we present next a simple mean for incorporating convex losses for $h_{i}(\cdot)$ at zero additional computational cost. To remind the reader, given a strictly convex function $\phi:\mathbb{R}\to\mathbb{R}$, the Bregman divergence between $x\in\mathbb{R}$ and $y\in\mathbb{R}$ is defined as, $D_{\phi}(x,y)=\phi(x)-(\phi(y)+\nabla\phi(y)(x-y))$. Our generalization amount to setting, $h_{i}(x)=D_{\phi}(x,y_{i})$ and keeping $g_{i}(\delta)$ intact, thus retaining for the fused lasso $g_{i}(\delta)=\lambda_{i}|\delta|$. This choice for $h_{i}$ implies that $\nabla h_{i}(x)=\nabla\phi(x)-\nabla\phi(y_{i})$. Since $\phi(\cdot)$ is strictly convex $\nabla\phi(\cdot)$ is continuous and strictly increasing. Following the same line of derivation as above we get that the recursive mappings for each $\mathsf{a}_{i}(\cdot)$ takes the following form,
$\mathsf{a}_{i}(\mathsf{x})=\Big{[}\mathsf{a}_{i+1}(\mathsf{x})+\nabla\phi(\mathsf{x})-\nabla\phi(y_{i+1})\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}~{}.$ (8)
This implies that $\mathsf{a}_{i}(.)$ is a piecewise linear as a function of $\nabla\phi(.)$. We thus follow the same procedure above whilst replacing $\mathsf{x}\mapsto\nabla\phi(x)$ for each linear segment in $\mathsf{a}_{i}$. This generalization to Bregman divergences is applicable throughout the remainder of the section. We keep using the squared error For brevity of our derivation in the sequel.
3.4 Isotonic Regression
In isotonic regression the goal is to find the best non-decreasing sequence, $x_{1}\leq x_{2}\leq\ldots\leq x_{i}\leq x_{i+1}\leq\ldots\leq x_{n}$ for which the sum of the losses $h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}$ is minimized. We replace the fused penalties with $g_{i}(\delta)=\mathbf{1}_{\delta\leq 0}$. This choice implies that $g_{i}^{*}(\alpha)=\mathbf{1}_{\alpha\geq 0}$ and its subgradient is,
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=0\\ 0&\alpha>0\end{cases}~{}.$
Parallel to the fused lasso, there exists a succinct representation for isotonic regression of the form $S_{i}=\{(\mathsf{x},\mathsf{a}_{i}(\mathsf{x})):\mathsf{x}\in\mathsf{R}\}$ which is prescribed recursively as,
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{+}$
where $[z]_{+}=\max(z,0)$. The derivation follows the same line as above as there exists a unique boundary value $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=0$ and the resulting description of $S_{i}$ is obtained from the union of,
$S_{i}\cap\{\mathsf{x}:\mathsf{x}>z_{i}^{-}\}$ $=$ $\{(\mathsf{x},\alpha_{i}^{\prime}(\mathsf{x})):\mathsf{x}>z_{i}^{-}\}$
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},0):\mathsf{x}\leq z_{i}^{-}\}~{}.$
We can in fact entertain a simple generalization which encompasses both absolute difference penalties, isotonic regression, and asymmetric difference penalties by defining,
$g_{i}(\delta)=\bar{\lambda}_{i}[\delta]_{+}\,-\,\underline{\lambda}_{i}[-\delta]_{+}\,\mbox{ with }\,\bar{\lambda}_{i}\geq 0\,,\,\underline{\lambda}_{i}\leq 0~{}.$
The recursive mapping takes the following form,
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{\underline{\lambda}_{i}}^{\bar{\lambda}_{i}}~{}.$
Setting $-\underline{\lambda}_{i}=\bar{\lambda}_{i}=\lambda_{i}$ yields the fused lasso whereas choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\infty$ gives isotonic regression. In addition, choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\lambda$ recovers the setting of [17]. To conclude this section we would like to note that isotonic regression can be used as the core for finding the closest vector to $\mathbf{y}$ subject to a $\leq k$-modality constraint. Due to submodularity of the objective the total runtime would be $\mathcal{O}(kn)$.
Figure 3: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for isotonic regression.
3.5 Fused Barriers
In certain data science applications the goal is to eliminate outlier points in an observed sequence rather than smooth or filter the sequence. We cast the problem as requiring the differences between consecutive points not to exceed prescribed maximal changes $b_{i}$ by establishing a series of barrier functions, $g_{i}(\delta)=\mathbf{1}_{|\delta|\leq\lambda_{i}}$. Each $b_{i}$ here can be thought of the maximum tolerance to changes and is not necessarily small and thus restrictive. For this choice of $g_{i}$ we get,
$\partial g_{i}^{*}(\alpha)=\begin{cases}-\lambda_{i}&\alpha<0\\ [-\lambda_{i},\lambda_{i}]&\alpha=0\\ \lambda_{i}&\alpha>0\end{cases}~{}.$ (9)
We next show that there exists a succinct representation of $S_{i}$ in the form of a piecewise linear map, $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$ albeit being slightly more involved. First, let us reuse the intermediate function $\mathsf{a}_{i}^{\prime}$ from the derivation of the fused lasso algorithm, $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. We denote by $\kappa_{\lambda}(\cdot)$ the Kvetsh operator (from Yiddish, literally “to squeeze, shrink”), colloquially called shrinkage,
$\kappa_{\lambda}(z)=\text{sign}(z)\,\big{[}|z|-\lambda\big{]}_{+}~{}.$
The construction of $\mathsf{a}_{i}$ from $\mathsf{a}_{i+1}$ consists of the following steps,
$\mathsf{a}_{i}^{\prime}(\mathsf{x})$ $:=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$ (Add Coordinate)
$z_{i}$ $:=z\,\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z)=0$ (Zero Crossing)
$\mathsf{a}_{i}(\mathsf{x})$ $:=\mathsf{a}_{i}^{\prime}\big{(}z_{i}+\kappa_{\lambda_{i}}(\mathsf{x}-z_{i})\big{)}$ (Kvetsh)
Informally, after constructing the intermediate function $\mathsf{a}^{\prime}_{i}$ which is strictly monotone, we find the point $z_{i}$ where $\mathsf{a}^{\prime}_{i}$ crosses zero. We then insert an interval of zero slope whose center is $z_{i}$ and spans from $z_{i}-\lambda_{i}$ through $z_{i}+\lambda_{i}$.
We now provide the details of the derivation of $\mathsf{a}_{i}$. Since $z_{i}$ is the locus of the zero of $\mathsf{a}^{\prime}_{i}$ then for $\mathsf{x}^{\prime}>z_{i}$ we have $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})>0\Rightarrow\partial g_{i}^{*}(\mathsf{a})=\lambda_{i}$ based on Eq. 9. Thus, for any $\mathsf{x}=\mathsf{x}^{\prime}+\lambda_{i}>z_{i}+\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\in S_{i}$. Analogously, for $\mathsf{x}<z_{i}-\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\in S_{i}$. Last, for $(z_{i},\mathsf{a}^{\prime}_{i}(z_{i}))=(z_{i},0)$ we have $\partial g_{i}^{*}(0)=[-\lambda_{i},\lambda_{i}]$. Therefore, for any $\mathsf{x}$ such that $z_{i}-\lambda_{i}\leq\mathsf{x}\leq z_{i}+\lambda_{i}$ it holds that $\mathsf{x}-z_{i}\in\partial g_{i}^{*}(0)\Rightarrow(\mathsf{x},0)\in S_{i}$. The resulting $S_{i}$ is obtained from the union of,
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$
[…]
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}~{}.$
Eq. (Kvetsh) provides a succinct functional form of the union of these three sets.
Since each $\mathsf{a}_{i}(\cdot)$ is a piecewise linear function we can use the same algorithmic skeleton used for instance in the fused lasso. Alas, in order to construct $\mathsf{a}_{i}$ we need to traverse $\mathsf{a}^{\prime}_{i}$ and find the locus, $z_{i}$, of its zero. This search would require $\mathcal{O}(n)$ time and thus the total run time would amount to $\mathcal{O}(n^{2})$. We provide in Appendix A a describe more efficient implementation that require a more elaborated data structure in order to reduce the amortized runtime to $\mathcal{O}(n\log(n))$.
Figure 4: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for fused barriers.
4 Variationally Penalized Multivariate Regression
We now extend the subgradient following framework to multivariate settings. For concreteness we examine objectives for instance $i$ of the form,
$h_{i}({\mathbf{x}})\,=\,\frac{1}{2}\big{\|}{\mathbf{x}}-\mathbf{y}_{i}\big{\|}^{2}_{{Q}_{i}}\,=\,\frac{1}{2}({\mathbf{x}}-\mathbf{y}_{i})^{\top}\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}~{}{\mathbf{x}},\mathbf{y}_{i}\in\mathbb{R}^{n}~{}.$ (10)
We impose the condition that the quadratic form is strictly positive definite, $\mathsf{Q}_{i}\succ 0$. The multivariate is intrinsically more complex than the scalar case due to the penalty entanglement of ${\mathbf{x}}_{i}$ with itself and with ${\mathbf{x}}_{i+1}$. To mitigate this difficulty we examine first the squared Euclidean distance between ${\mathbf{x}}_{i}$ and ${\mathbf{x}}_{i+1}$,
$g_{i}(\boldsymbol{\delta})\,=\,\frac{1}{2}\lVert{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\rVert^{2}\,=\,\frac{1}{2}\lVert\boldsymbol{\delta}\rVert^{2}~{}.$
This setting generalizes squared fused penalty for the scalar case for which $Q_{i}=[q_{i}\in\mathbb{R}]$ and thus can be absorbed into $\lambda_{i}$. For this setting we have,
$\nabla h_{i}({\mathbf{x}})=\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}\mbox{and}~{}\nabla g_{i}^{*}(\boldsymbol{\delta})=\boldsymbol{\delta}~{}.$
The derivation of subgradient following for this setting also serves as a building block in the more complex multivariate settings presented in the sequel.
The Fenchel dual in the multivariate case where $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined as,
$f^{*}(\boldsymbol{\alpha})=\sup_{{\mathbf{x}}\in\mathrm{dom}(f)}\big{\{}\boldsymbol{\alpha}^{\top}{\mathbf{x}}-f({\mathbf{x}})\big{\}}~{}.$
Since both $h_{i}(\cdot)$ and $g_{i}(\cdot)$ defined above are strictly convex, as we require $\mathsf{Q}_{i}\succ 0$ the Fenchel dual for all $i\in[n]$ are unique and satisfy for all feasible dual pairs $({\mathbf{x}},\boldsymbol{\alpha})$ the equation $\boldsymbol{\alpha}=\nabla\!f({\mathbf{x}})$ typically referred to as the link function or the mirror map. We use this property in the ensuing derivation.
Since the gradients’ components are of unary form and linear the subgradient sets can be written succinctly as,
$S_{i}=\left\{({\mathbf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}{\mathbf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+{\mathbf{c}}_{i}=\mathbf{0}\right\}~{}.$
Since the mirror map here is bijective the admissible sets of (6) and (6) can be written as,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!\tilde{f}_{i+1}(\boldsymbol{\mathsf{x}}_{i+1})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!f_{i}(\boldsymbol{\mathsf{x}}_{i})\big{\}}~{}.$
We therefore can easily unravel the recursive form for $\mathsf{A}_{i-1}$, $\mathsf{B}_{i-1}$, ${\mathbf{c}}_{i-1}$ from $\mathsf{A}_{i}$, $\mathsf{B}_{i}$, ${\mathbf{c}}_{i}$,
$S_{i-\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{a}}-\boldsymbol{\mathsf{a}}^{\prime}=\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i});\;(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}^{\prime})\in S_{i}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\left(\boldsymbol{\mathsf{a}}-\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i})\right)+{\mathbf{c}}_{i}=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{},$ $\vskip 3.0pt plus 1.0pt minus 1.0ptS_{i-1}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{x}}^{\prime}=\boldsymbol{\mathsf{a}};\;(\boldsymbol{\mathsf{x}}^{\prime},\boldsymbol{\mathsf{a}})\in S_{i-\frac{1}{2}}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})(\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{a}})+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+(\mathsf{B}_{i}(\mathsf{I}+\mathsf{Q}_{i})-\mathsf{A}_{i})\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{}.$
We therefore get that,
$\mathsf{A}_{i-1}$ $=$ $\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i}$
$\mathsf{B}_{i-1}$ $=$ $\mathsf{B}_{i}(\mathsf{I}+\mathsf{Q}_{i})-\mathsf{A}_{i}$
${\mathbf{c}}_{i-1}$ $=$ ${\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i}~{}.$
For boundary conditions we require,
$(\mathsf{A}_{n},\mathsf{B}_{n},{\mathbf{c}}_{n})=(\mathsf{O},\,\mathsf{I},\,\mathbf{0})~{}\mbox{ and }~{}(\mathsf{A}_{1}-\mathsf{B}_{1}\mathsf{Q}_{1}){\mathbf{x}}_{1}+({\mathbf{c}}_{1}+\mathsf{B}_{1}\mathsf{Q}_{1}\mathbf{y}_{1})=0~{}.$
Upon obtaining ${\mathbf{x}}_{1}$ along with $\boldsymbol{\alpha}_{0}\equiv\mathbf{0}$, we construct the trajectory of the solution by following a backward recursion,
$\boldsymbol{\alpha}_{i}$ $=$ $\boldsymbol{\alpha}_{i-1}-\mathsf{Q}_{i}({\mathbf{x}}_{i}-\mathbf{y}_{i})$
${\mathbf{x}}_{i+1}$ $=$ ${\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}~{}.$
5 Sparse Variationally Penalized Multivariate Regression
We use ${\mathbf{x}}_{1:n}$ henceforth to denote the sequence of vectors ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$. While the algorithm of Sec. 4 entertains a concise form it does not yield variational sparsity. To do so we can readily use the $1$-norm for the variational penalty. This would yield though element-wise fusion whereas full variational sparsity entails that for some indices $i\in[n]$, ${\mathbf{x}}_{i}={\mathbf{x}}_{i+1}$. In this section we use the algorithm from as a building block while replace the norm squared of ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ with either the $2$-norm of the difference vector or its infinity norm. For brevity, we assume throughout this section that $\forall i,\lambda_{i}=1$ and are thus omitted.
5.1 Multivariate Regression with $\ell_{2}$-Variational Penalty
We first replace each of the norm-squared variational penalties with a norm, thus the goal is to find the minimizer of,
$\phi({\mathbf{x}}_{1:n})=\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|~{}~{}.$ (11)
Alas, this formulation does not entail a closed form subgradient following procedure. The approach that we take here is to construct a relaxation that yields the follow convergent algorithm.
Algorithm 1 Surrogate Multivariate Nariat
  initialize ${\mathbf{x}}_{1:n}^{0}=\mathbf{y}_{1:n}$
  for $t=1\ \textbf{to}\ t_{\text{max}}$ do
     $\forall i:\rho_{i}^{t}=\max\big{(}\|{\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t}\|,\,\epsilon\big{)}$
     ${\mathbf{x}}^{t+1}_{1:n}=\displaystyle\operatorname*{\arg\!\min}_{{\mathbf{x}}_{1:n}}\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\frac{\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}}{2\rho_{i}^{t}}$
  end for
  return ${\mathbf{x}}^{t_{\text{max}}}_{1:n}$
We next show that Algorithm 1 converges linearly to the minimum of,
$\hat{\phi}({\mathbf{x}}_{1:n})=\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})~{}~{}.$ (12)
where $g_{\epsilon}(\cdot)$ is the Huber [10] loss defined as,
$g_{\epsilon}(\boldsymbol{\delta})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}\|\leq\epsilon\\ \|\boldsymbol{\delta}\|-\frac{\epsilon}{2}&\|\boldsymbol{\delta}\|>\epsilon\end{cases}~{}.$
Since for any $\boldsymbol{\delta}$ it holds that $\|\boldsymbol{\delta}\|-\frac{\epsilon}{2}\leq g_{\epsilon}(\boldsymbol{\delta})\leq\|\boldsymbol{\delta}\|$, the solution of the relaxed problem Eq. (12) is at most $\frac{(n-1)\epsilon}{2}$ suboptimal with respect to the original problem defined in Eq. (11). Let us introduce a further auxiliary function,
$\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}^{\prime}\|\leq\epsilon\\ \frac{\|\boldsymbol{\delta}\|^{2}}{2\|\boldsymbol{\delta}^{\prime}\|}+\frac{\|\boldsymbol{\delta}^{\prime}\|-\epsilon}{2}&\|\boldsymbol{\delta}^{\prime}\|>\epsilon\end{cases}~{}.$
The function $\hat{g}_{\epsilon}$ is a quadratic upper bound of $g_{\epsilon}(\boldsymbol{\delta})$ expanded at $\boldsymbol{\delta}^{\prime}$,
$\forall\boldsymbol{\delta},\boldsymbol{\delta}^{\prime}:~{}\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})\geq g_{\epsilon}(\boldsymbol{\delta})~{}\mbox{ and }~{}\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta})=g_{\epsilon}(\boldsymbol{\delta})~{}.$
By carefully examining each cases we have the following,
$\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})$ $=$ $g_{\epsilon}(\boldsymbol{\delta}^{\prime})+\nabla{g_{\epsilon}}(\boldsymbol{\delta}^{\prime})^{\top}(\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime})+\frac{1}{2\,\max\big{(}\|\boldsymbol{\delta}^{\prime}\|,\epsilon\big{)}}\|\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime}\|^{2}$
$\leq$ $g_{\epsilon}(\boldsymbol{\delta})+\frac{1}{2\,\max\big{(}\|\boldsymbol{\delta}^{\prime}\|,\epsilon\big{)}}\|\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime}\|^{2}~{}.$
Hence, Algorithm 1 is a proximal method where we iteratively apply the multivariate subgradient following from previous section to obtain an exact solution for each adaptively constructed upper bound. We follow a similar skeletal proof to show the convergence.
We denote by $\eta^{t}=\min_{i}\rho_{i}^{t}\geq\epsilon$ and the optimum of the relaxed problem as $\hat{\phi}^{*}=\hat{\phi}({\mathbf{x}}_{1:n}^{*})$. We next prove a sequence of bounds as the means to show convergence starting with employing the definition of $g_{\epsilon}$ and $\hat{g}_{\epsilon}$,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}^{t+1}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}^{t+1}-{\mathbf{x}}_{i}^{t+1})$
$\leq$ $\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}^{t+1}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\hat{g}_{\epsilon}({\mathbf{x}}_{i+1}^{t+1}-{\mathbf{x}}_{i}^{t+1},{\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t})$
$=$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\hat{g}_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i},{\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t})$
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\frac{1}{2\rho_{i}^{t}}\|({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-({\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t})\|^{2}$
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\frac{1}{\rho_{i}^{t}}\big{(}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i+1}^{t}\|^{2}+\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\|^{2}\big{)}$
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\sum_{i=1}^{n}\frac{2}{\eta^{t}}\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\|^{2}~{}.$
Next we introduce a dummy sequence ${\mathbf{z}}_{1:n}(\mu)=(1-\mu)\,{\mathbf{x}}_{1:n}^{t}+\mu\,{\mathbf{x}}_{1:n}^{*}$ and upper bound the last term by a minimizer constrained to the segment from ${\mathbf{x}}_{1:n}^{t}$ to ${\mathbf{x}}_{1:n}^{*}$ and then use the convexity of the objective,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$$\leq$ $\min_{\mu\in[0,1]}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{z}}_{i}(\mu)-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{z}}_{i+1}(\mu)-{\mathbf{z}}_{i}(\mu))+\sum_{i=1}^{n}\frac{2}{\eta^{t}}\|{\mathbf{z}}_{i}(\mu)-{\mathbf{x}}_{i}^{t}\|^{2}$
$\leq$ $\min_{\mu\in[0,1]}(1-\mu)\,\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\mu\,\hat{\phi}({\mathbf{x}}_{1:n}^{*})+\sum_{i=1}^{n}\Big{(}\frac{2\mu^{2}}{\eta^{t}}-\frac{\mu(1-\mu)}{2}\Big{)}\|{\mathbf{x}}_{i}^{t}-{\mathbf{x}}_{i}^{*}\|^{2}~{}.$
Last, we replace the minimizer w.r.t. $\mu$ with a specific choice $\mu=\frac{\eta^{t}}{4+\eta^{t}}$ and get,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$$\leq$ $\frac{4}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\frac{\eta^{t}}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{*})~{}.$
Therefore, the optimality gap of the surrogate $\hat{\phi}$ satisfies,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})-\hat{\phi}^{*}\leq\frac{4}{4+\eta^{t}}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}\leq\frac{4}{4+\epsilon}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}~{}.$
Thus,we get that the sequence $\hat{\phi}({\mathbf{x}}_{1:n}^{t})$ convergence linearly to $\hat{\phi}^{*}$.
To conclude the section would like to underscore the advantage of the combination of using a surrogate loss with exact subgradient following over a dual ascent algorithm tailored for the problem. The dual proximal gradient (DPG) is described and briefly analyzed in Appendix B. To do so, we generated synthetic data of varying levels of variational sparsity as follows. For each sparsity level, let $\mathcal{D}$ denote a distribution whose density function statisfies $\mathcal{D}({\mathbf{x}})\sim\text{e}^{-\lambda\|{\mathbf{x}}\|}$ such that its probability mass within the unit ball is equal to a prescribed fusion probability. We first sample a sequence of Boolean variables $\{\nu_{i}\}$ from the Bernoulli distribution with a predefined probability which designated the chance of fusing two consecutive vectors. When $\nu_{i}=1$, we sampled dual variables $\boldsymbol{\alpha}_{i}$ from $\mathcal{D}$ over unit ball and set ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}=\mathbf{0}$. Otherwise, we sampled ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ from $\mathcal{D}$ outside the unit ball and set $\boldsymbol{\alpha}_{i}\in\partial\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. We then constructed $\mathbf{y}_{i}$ based on the optimality condition, $\mathbf{y}_{i}={\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}+\boldsymbol{\alpha}_{i-1}$. This randomized generation process ensures that ${\mathbf{x}}_{1:n},\boldsymbol{\alpha}_{0:n}$ are primal-dual optimal for $\mathbf{y}_{1:n}$.
We run both DPG and Nariat for 100 iterations each and then merged two consecutive points when the $2$-norm of the variational difference was below a threshold of $10^{-8}$. For each level of variational sparsity we generated $100$ random sequences of vectors with $d=100$ and $n=100$. We measured the $2$-norm between the solution found by the algorithms and the primal-dual optimal solution which was used to generate the synthetic data. The results are shown in Figure 5. The results clearly indicate that Nariat is superior to DPG which was tailored for the problem. The gap between ${\mathbf{x}}^{*}_{1:n}$ and $\hat{{\mathbf{x}}}_{1:n}$ found by Nariat is insensitive of the fusion probability whereas DPG’s performance deteriorates as the level of sparsity grows. Equally if not more important, Nariat exhibits much superior recovery rate of fusion events and in fact seems to slightly improve with the level of sparsity. In contrast, DPG’s rate of recovery rapidly falls to the point that at a sparsity level of $70\%$ almost no fusion events were identified.
Figure 5: Nariat vs. DPG for sparse multivariate discovery using the the $2$-norm .
5.2 Multivariate Regression with $\infty$ norm as variational Penalty
The second vector fusion penalty we consider is the infinity norm of the difference vectors ${\mathbf{x}}_{i}-{\mathbf{x}}_{i+1}$ which entails the problem,
$\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}~{}~{}.$ (13)
We associate a slack variable with each difference vector $\xi_{i}$ and rewrite the problem as,
$\min_{{\mathbf{x}}_{1:n},\boldsymbol{\xi}\geq 0}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}.$
We cast the problem as a gradient-based search for the optimum $\xi_{1:n-1}^{*}$. On each gradient step we solve a residual problem in ${\mathbf{x}}$ using the subgradient following procedure with bounded variation from the previous section. Formally, given any $\xi_{1:n-1}\geq 0$ we use the problem given by Eq. 13 to define $\mathcal{L}(\xi_{1:n-1})$ as follows,
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}.$
In order to compute the subgradient of $\mathcal{L}$, we introduce the dual variables and write the constrained problem as,
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\,\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}+\sum_{i=1}^{n-1}\boldsymbol{\alpha}_{i}^{\top}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-\xi_{i}\|\boldsymbol{\alpha}_{i}\|_{1}$
Due to strong duality we can interchange min and max in the minimax problem to get,
$\mathcal{L}(\xi_{1:n-1})$ $=$ $\max_{\boldsymbol{\alpha}_{1:n-1}}\,\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}+\sum_{i=1}^{n-1}\boldsymbol{\alpha}_{i}^{\top}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-\xi_{i}\|\boldsymbol{\alpha}_{i}\|_{1}$
$=$ $\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\|_{2}^{2}+\sum_{i=1}^{n-1}\xi_{i}\big{(}1-\|\boldsymbol{\alpha}_{i}\|_{1}\big{)}~{}.$
Denote by $\boldsymbol{\alpha}_{1:n-1}^{*}$ the maximizer of the above equation and let $\mathbf{u}$ be the vector whose $i$’th coordinate is $u_{i}=1-\|\boldsymbol{\alpha}_{i}^{*}\|_{1}$. For any $\xi_{1:n-1}^{\prime}\geq 0$, we have
$\mathcal{L}(\xi_{1:n-1})+\mathbf{u}^{\top}(\xi_{1:n-1}^{\prime}-\xi_{1:n-1})=$
$\displaystyle\frac{1}{2}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}^{*}-\boldsymbol{\alpha}_{i-1}^{*}\|_{2}^{2}+\sum_{i=1}^{n-1}\xi_{i}^{\prime}\big{(}1-\|\boldsymbol{\alpha}_{i}^{*}\|_{1}\big{)}\leq\mathcal{L}(\xi_{1:n-1}^{\prime})~{}.$
Therefore $\mathbf{u}\in\partial\mathcal{L}(\xi_{1:n})$, namely, it is a subgradient of $\mathcal{L}$ at $\xi_{1:n-1}$. Again by strong duality, we can obtain the optimal dual variables by first solving the primal problem and then infer the dual variables $\boldsymbol{\alpha}_{:}$ from optimality conditions. Given $\xi_{1:n-1}$ the primal is separable and we thus can apply the subgradient following from 3.5 to obtain both the optimal primal and dual variables. The slack variables $\xi_{1:n-1}$ are updated using gradient stepping. To do so, we derive upper bounds such that $\xi_{i}^{*}\leq D_{i}$ and then apply AdaGrad [5] with an initial per-coordinate learning rate of $D_{i}/\sqrt{2}$. The derivation of the upper bounds is given in Appendinx C.
[…]
Figure 6: Nariat vs. DPG for sparse multivariate discovery using the the $\infty$-norm .
6 High-Order Variational Penalties
The differences $x_{i+1}-x_{i}$ can morally be viewed as discrete derivatives. The focus of this section is higher order discrete derivatives as variational penalties. For example, penalizing for the second order derivative amounts to,
$\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\lambda_{i}\big{|}(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})\big{|}~{}~{}.$
Rather than discussing specific forms of discrete difference operators, we analyze a general case stemming from convolution operators. Formally, let $\boldsymbol{\tau}=(-1,\boldsymbol{\tilde{\tau}})$ where $\boldsymbol{\tilde{\tau}}=(\tau_{1},\tau_{2},\ldots,\tau_{k})$ be vectors over $\mathbb{R}^{k+1}$ and $\mathbb{R}^{k}$ respectively. The convolution of ${\mathbf{x}}=(x_{1},\ldots,{\mathbf{x}}_{n})$ with $\boldsymbol{\tau}$ is a vector whose $i$’th entry is, $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=\sum_{m=0}^{k}\tau_{m}\,x_{i-m}~{}.$ Equipped with the aforementioned definitions we now focus on problems of the following general form,
$\min_{\mathbf{x}}\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=k}^{n-1}g_{i}\big{(}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{)}~{}.$ (14)
Note that we can choose $\tau_{0}=-1$ as scaling is a degree of freedom that can be absorbed into $g_{i}$. This constructions implies that $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=(\boldsymbol{\tilde{\tau}}\ast{\mathbf{x}})_{i}-x_{i}$. Here we on purpose shift the indices of $g_{i}$ by $1$ to keep the same notation as used in the algorithms above when plugging in $\boldsymbol{\tilde{\tau}}=(1)~{}.$
We use $\boldsymbol{\tau}\circledast{\mathbf{v}}$ to denote the convolution of a vector ${\mathbf{v}}$ with the reversal of $\boldsymbol{\tau}$ which we define as,
$(\boldsymbol{\tau}\circledast{\mathbf{v}})_{i}=\sum_{m=0}^{k}\tau_{m}v_{i+m}~{}.$
We again confine ourselves to settings for which strong duality holds thus,
$U$ $=$ $\min_{{\mathbf{x}}_{1:k}}\,\max_{\alpha_{k}}\,\min_{x_{k+1}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=k}^{n-1}\left[\alpha_{i}(\boldsymbol{\tau}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})\right]$
Analogously, we define a sequence of intermediate functions defined as
$f_{i}({\mathbf{x}})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(\boldsymbol{\tau}*{\mathbf{x}})_{j+1}-g_{j}^{*}(\alpha_{j})\right]~{}.$ (15)
This definition means that $f_{i}:\mathbb{R}^{i}\to\mathbb{R}$ and thus formally defined as $f_{i}({\mathbf{x}}_{1:i})$ since ${\mathbf{x}}_{i+1:n}$ is grounded to its optimal vector. As boundary conditions we set $f_{n}({\mathbf{x}})=0$. Similar to the constructions above we introduce an auxiliary function, $\tilde{f}_{i}({\mathbf{x}})=f_{i}({\mathbf{x}})+h_{i}(x_{i})$ which, like $f_{i}$, is a function of ${\mathbf{x}}_{1:i}$
To simplify notation we denote by $\boldsymbol{\mathsf{x}}_{i}$ the prefix ${\mathbf{x}}_{1:i}$ and $\boldsymbol{\mathsf{a}}_{i}$ for $\boldsymbol{\alpha}_{i:n-1}$. Albeit the analogous notation, $\boldsymbol{\mathsf{x}}_{i}$ denotes a free variable whereas $\boldsymbol{\mathsf{a}}_{i}$ designates the optimal dual vector. Using these definitions we introduce a summation-free form for $f_{i}$,
$f_{i}({\mathbf{x}})$ $=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\tilde{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+1}({\mathbf{x}})+h_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$ (16)
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\tilde{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{\tilde{f}_{i+1}(x_{i+1}|\boldsymbol{\mathsf{x}}_{i})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\tilde{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})-\tilde{f}_{i+1}^{*}(\alpha_{i}|\boldsymbol{\mathsf{x}}_{i})~{},$
where $\tilde{f}_{i+1}(\cdot|\boldsymbol{\mathsf{x}}_{i})=\tilde{f}_{i+1}(\boldsymbol{\mathsf{x}}_{i},\cdot)$.
The subgradient conditions for locally-restricted optimality define two admissible sets,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial\tilde{f}_{i+1}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}~{}.$
We next show that the following recursive form holds,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\tilde{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}.$
Notice that the admissible set consists of the full sequence of locally-admissible pairs. Due to strong duality, $S_{i}$ actually defines the point-to-set mapping from a prefix vector $\boldsymbol{\mathsf{x}}_{i}$ to its dual solutions of $f_{i}(\cdot)$. Hence, we can rewrite 15 as,
$f_{i}({\mathbf{x}})$ $=$ $\max_{\boldsymbol{\alpha}_{i:n-1}}\,\min_{{\mathbf{x}}_{i+1:n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(\boldsymbol{\tau}*{\mathbf{x}})_{j+1}-g_{j}^{*}(\alpha_{j})\right]$
$=$ $\max_{\boldsymbol{\alpha}_{i:n-1}}\>\sum_{j=i}^{n-1}\left[\alpha_{j}(\boldsymbol{\tau}*\boldsymbol{\mathsf{x}}_{i})_{j+1}-h_{j+1}^{*}(-(\boldsymbol{\tau}\circledast\boldsymbol{\alpha})_{j})-g_{j}^{*}(\alpha_{j})\right]~{}.$
Thus, given any $\boldsymbol{\mathsf{x}}_{i}$ and for any $\boldsymbol{\mathsf{a}}_{i}$ attaining the above maximum we have
$\forall m\in\{0,\ldots,k-1\}:\;(\boldsymbol{\tilde{\tau}}\circledast\boldsymbol{\mathsf{a}}_{i})_{-m}\in\partial_{x_{i-m}}f_{i}(\boldsymbol{\mathsf{x}}_{i})~{}.$
Now, given $S_{i+1}$ we conduct an expansion,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial f_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})+\partial h_{i+1}(x_{i+1})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\tilde{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(x_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}~{}.$
Using the fact that $\mathsf{a}_{i}\in\partial\tilde{f}_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial\tilde{f}^{*}_{i+1}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})$ we get,
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial\tilde{f}_{i+1}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\tilde{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}.$
Finally, to satisfy the boundary conditions we need to find the solution of,
$\min_{{\mathbf{x}}_{1:k}}f_{k}({\mathbf{x}})+\sum_{i=1}^{k}h_{i}(x_{i})~{}.$
The subgradient conditions imply that
[…]
where we used the above property to compute $\partial_{\mathsf{x}_{i}}f_{k}({\mathbf{x}})$.
Figure 7: Exact vs. approximate subgradient following for second-order variational penalties.
Second-order Variational Penalties.
To illustrate the power of the convolution-based representaion, we provide a specific derivation for variational penalties which correspond to second order finite differences. To do so we set $\boldsymbol{\tilde{\tau}}=(2,-1)$. Since $f_{i}(\cdot)$ depends only on $x_{i-1}$ and $x_{i}$, we can represent the subgradient sets as,
$S_{i}=\{\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})\;;\;\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})\}~{}\mbox{ and }~{}S_{i+\frac{1}{2}}=\{\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}.$
Assume this holds true for $S_{i+1}~{}\mbox{ i.e. }~{}S_{i+1}=\{\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}.$ The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})=2\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})-\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})+\mathsf{x}_{i+1}-y_{i+1}~{}\mbox{ and }~{}\mathsf{a}_{i+1}^{\prime}(\cdot)=\mathsf{a}_{i+1}(\cdot)~{}.$
To obtain the recursive map $S_{i+\frac{1}{2}}\to S_{i}$, we need to solve for $\mathsf{x}_{i+1}$ from the following equation
$2\mathsf{x}_{i}-\mathsf{x}_{i-1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1}))~{}.$
Similar to the fused lasso, the solution is given by the backward map $T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})$ defined as
$T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\big{[}2\mathsf{x}_{i}-\mathsf{x}_{i-1}\big{]}_{z_{i+1}^{-}(\mathsf{x}_{i})}^{z_{i+1}^{+}(\mathsf{x}_{i})}~{},$
where $z_{i+1}^{\pm}(\mathsf{x}_{i})$ denotes the boundary values such that $\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},z_{i+1}^{\pm}(\mathsf{x}_{i}))=\pm 1~{}.$ Using this form of $T$ back in $\mathsf{x}_{i+1}$, we obtain
$\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))$
and
$\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))~{}.$
Here, $\mathsf{a}_{i}(\cdot)$ is a piecewise linear manifold in $\mathbb{R}^{2}$. We store triplets of the form $\{v_{j}^{1},v_{j}^{2},\alpha_{j}\}$ for vertices of each facets defining $\mathsf{a}_{i}$. We conjecture that the number of facets grows as $\mathcal{O}(n^{2})$ but leave formal analysis of runtime as future research. We empirically observed that the description of $\mathsf{a}_{i}$ is typically linear in the length of the input sequence. Alas, a disadvantage of the exact subgradient following described here is the need to maintain a data structure for high dimensional objects. We describe an alternative approximation yielding a simple iterative procedure for the problem in Appendix D.
We next demonstrate the effectiveness of exact subgradient following over the approximation-based method. To do so, we generated synthetic data of varying sequence length similar to previous section. We sampled $(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})$ from $\mathcal{D}$ and then constructed $\mathbf{y}$ from optimality conditions. Throughout the experiment, the sparsity level was set to $50\%$. In each experiment, we ran the approximation-based algorithm for $100$ iterations. The results are summarized in Figure 7. The advantage of exact subgradient following over approximation both in terms of runtime and optimality is apparent. The exact subgradient following depends rather weakly on the problem size whereas the quality of the solution as well run time quickly deteriorates even for short sequences.
7 Sub-Discussion
There are several possible extensions of the framework presented in this paper. The univariate penalties are readily extendable to non-symmetric counterparts and the addition of insensitivity regions. For instance, a generalization the symmetric penalty of the fused lasso amounts to,
$\lambda_{i}^{\mathsf{R}}\big{[}x_{i+1}-x_{i}-\epsilon_{i}^{\mathsf{R}}\big{]}_{+}\,+\,\lambda_{i}^{\mathsf{L}}\big{[}x_{i}-x_{i+1}-\epsilon_{i}^{\mathsf{L}}\big{]}_{+}~{}.$
Bu choosing $\lambda_{i}^{\mathsf{R}}=\lambda_{i}^{\mathsf{L}}=\lambda_{i}$ and $\epsilon_{i}^{\mathsf{R}}=\epsilon_{i}^{\mathsf{L}}=0$ we obtain the fused lasso’s penalty. To encompass this version we simply need to compute $\partial g_{i}^{*}$ from which we derive a new mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. The subgradient procedure uses a similar form for the mapping given in Eq. 7 though the introduction of insensitivity regions to the left and/or the right of the loci where $x_{i}=x_{i+1}$ would introduce two addition inflection points. The resulting procedure for constructing the subgradient would require at most $\mathcal{O}(n\log n)$ time.
We limited our discussion to sequential penalties which can described as a chain of local connectivities $x_{1}\leftrightarrow x_{2}\leftrightarrow\ldots x_{i}\leftrightarrow x_{i+1}\ldots$ forming a left-to-right dependency graph. Kolmogorov et. al [12] examined a setting where the total variation penalties form a tree. Let us denote the variation dependency graph by $G=(V,E)$ where $V=[n]$ and $(i,j)\in E$ if there exists a penalty term for the variables $x_{i}$ and $x_{j}$. That is, the penalized problem takes the form of,
$U=\min_{{\mathbf{x}}}\>\sum_{i}\frac{1}{2}(x_{i}-y_{i})^{2}+\sum_{(i,j)\in E}|x_{i}-x_{j}|~{}~{}.$
In settings for which strong duality holds the above problem can be re-parameterized as,
$U$ $=$ $\max_{\|\boldsymbol{\alpha}\|_{\infty}\leq 1}\,\min_{{\mathbf{x}}}\>\frac{1}{2}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}+\sum_{(i,j)\in E}\alpha_{i,j}(x_{i}-x_{j})~{}.$
When $E$ defines a tree the subgradient following procedure can utilize the form described in [12] albeit increasing the time complexity to $\mathcal{O}(n^{2})$.
[Uncaptioned image][Uncaptioned image]
This work opens numerous directions for applied research. We would like here to sketch a single domain of potential applications of trajectory smoothing. Consider the task of mimicking cursively written notes such as the word foundation depicted on the top right. While the writing is clearly intelligible the letters are not of the same size and the word’s baseline drifts up and down. This characteristics is common to all cursively written texts by humans, and by some macaque monkeys. However, a synthetically generated text by a brain-controlled prosthetic would benefit from a low-jerk and smooth motor control of the electromechanical apparatus in charge of moving the artificial limb. To this end, a dynamical system such as the spring model of Hollerbach [9] can be used to approximate the noisy observations while inflicting penalty on high-jerk motions using the tools introduced in Sec. 6. The result would resemble the text depicted on the bottom right which was generated by a context-dependent dynamic programming.
Acknowledgments
We would like to thanks Tomer Koren and Yishay Mansour for fruitful discussions. The short discussion above is motivated by conversations with the late Prof. Tali Tishby who is missed dearly. This manuscript was written without the usage of AIML tools.
Appendix A Improving Runtime for Fused Barriers using Skiplist
We store each $\mathsf{a}_{i}(\cdot)$ as a linked-list data structure of implicit change points (vertices) and explicit linear intervals connecting two consecutive vertices. A vertex $(v_{j},\mathsf{a}_{i}(v_{j}))=(v_{j},\alpha_{j})$ represents a point where $\mathsf{a}_{i}(\cdot)$ changes slope. Edge $(d_{j},s_{j})$ stores the length and slope of the line segment between $v_{j}$ and $v_{j+1}$,
$d_{j}=v_{j+1}-v_{j}\,;\,s_{j}=(\alpha_{j+1}-\alpha_{j})/d_{j}~{}.$
This representation allows lazy evaluation in a similar fashion to the fused lasso. The Kvetsh step does not change the lengths of intervals except for the new segment encapsulating the zero crossing. Since the length of all other line segments remain intact this edge-based representation facilitates a stepping stone for an efficient implementation.
By construction of $\mathsf{a}_{i+1}$ we know that $\exists j$ for which $v_{j}=z_{i+1}-\lambda_{i+1}$ and $\mathsf{a}_{i+1}(z_{i+1}-\lambda_{i+1})=0$. This implies that in an explicit representation of $\mathsf{a}^{\prime}_{i}(\cdot)$ we would need to update vertex $j$, $(z_{i+1}-\lambda_{i+1},0)\mapsto(z_{i+1}-\lambda_{i+1},z_{i+1}-\lambda_{i+1}-y_{i+1})$. In the explicit yet lazy representation of the ensuing edge, the slope $s_{j}$ of the segment $[v_{j},v_{j+1}]$ becomes $s_{j}+n-i$. Next we find the zero crossing point of $\mathsf{a}^{\prime}_{i}(\cdot)$ by traversing the list of edges in the direction towards zero and locating the index $k$ for which $\alpha_{k}\leq 0$ and $\alpha_{k+1}>0$. In case such an index does not exist the zero of $\mathsf{a}^{\prime}_{i}(\cdot)$ is right of its right-most vertex. Once we identify the encapsulating interval, we calculate $z_{i}$ by solving a linear equation from $(v_{k},\alpha_{k})$ and $(v_{k+1},\alpha_{k+1})$. We next Kvetsh the list by inserting two vertices between $k$ and $k+1$ whose content is $(z_{i}-\lambda_{i},0)$ and $(z_{i}+\lambda_{i},0)$ and updating $d_{k}$ accordingly. This sequential traversal of the list in the worst-case would require linear time and thus total run time would remain intact at $\mathcal{O}(n^{2})$.
In order to facilitate a more efficient procedure for finding zero crossing. We make use of a skip-list data structure [14]. A skip-list is an extension of linked list with multiple layers and with which the expected search time is $\mathcal{O}(\log{n})$. The skip-list has ideally in our setting $\mathcal{O}(\log{n})$ layers where layer $k$ contains vertices whose index modulo $2^{k}$ is zero. The search procedure starts from highest layer and goes down once we reach the interval encapsulating zero. In practice, a skip-list is a randomized data structure where each vertex appears in a higher layer with probability $\frac{1}{2}$ to support insertion efficiently.
We follow the same lazy representation as above. We store $(d_{j}^{k},s_{j}^{k})=(v_{j+1}^{k}-v_{j}^{k},s_{j}^{k})$ for edge $j$ at layer $k$. We also add a precursor vertex at the beginning of each layer. This vertex contains the smallest possible value of $\{v_{j}\}$ as serves as a boundary point. To construct $\mathsf{a}^{\prime}_{i}(\cdot)$ we update $\alpha_{0}\leftarrow\alpha_{0}+v_{0}-y_{i+1}$. Then, we follow the search procedure above until reaching the lowest layer where we calculate $z_{i}$ by solving the same linear equation. During insertion of vertex $(z_{i}-\lambda_{i},0)$, we flip a coin to decide whether it should appear in a higher layer. Note that along the searching path, we have calculated $(v_{j}^{k},\alpha_{j}^{k})$ and $(v_{j+1}^{k},\alpha_{j+1}^{k})$ per layer which contains the locus of the zero crossing. Thus, whenever a vertex should appear in a prior (higher) layer, we use the corresponding values for updating the connecting edges. We perform an equivalent update for $(z_{i}+\lambda_{i},0)$. Last, we update the $v_{0}\leftarrow v_{0}-\lambda_{i}$. The end result is an $\mathcal{O}(n\log{n})$ amortized time algorithm.
Appendix B Dual Proximal Gradient
We give here further details on the dual proximal gradient method (DPG) we used in comparing our core algorithmic contribution. DPG merits further investigation on its own and to the best of our knowledge the short analysis with general norm penalties below was not derived in previous research.
[…]
$\min_{{\mathbf{x}}_{1:n}}~{}\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|_{2}^{2}+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|~{}.$
The dual of this problem is,
$\max_{\boldsymbol{\alpha}_{0:n}}~{}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\|_{2}^{2}~{}\;\text{s.t.}\;~{}\|\boldsymbol{\alpha}_{i}\|_{*}\leq 1~{}.$
For boundary conditions we set $\boldsymbol{\alpha}_{0}\equiv\boldsymbol{\alpha}_{n}\equiv\mathbf{0}$. Denote by $B_{*}=\{{\mathbf{x}}:\|{\mathbf{x}}\|_{*}\leq 1\}$ the unit ball w.r.t the dual norm and $\Pi_{B_{*}}(\cdot)$ be the projection operator onto the unit ball. In each iteration of dual proximal gradient,
$\boldsymbol{\alpha}_{i}\,\leftarrow\,\Pi_{B_{*}}\big{(}\boldsymbol{\alpha}_{i}-\eta(\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}+\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i+1}-\mathbf{y}_{i+1})\big{)}~{}.$
In order to underscore the advantage of the sub-gradient following method, let us briefly examine the condition number of the dual problem. Denote by $\mathsf{D}$ the difference matrix
$\mathsf{D}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\end{bmatrix}\in\mathsf{R}^{n-1\,\times\,n}~{}.$
The condition number of the dual problem is the same as that of $\mathsf{D}\mathsf{D}^{\top}$. If we choose $\boldsymbol{\alpha}$ such that
$\alpha_{i}=\frac{{(-1)}^{i}}{\sqrt{n-1}}~{},$
we have $\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=4-\frac{2}{n-1}$. Assume without loss of generality that $n$ is even and let us choose next
$\alpha_{i}=\frac{\frac{n}{2}-|i-\frac{n}{2}|}{\sqrt{n^{3}/12}}~{},$
for which we have $\|\boldsymbol{\alpha}\|_{2}\geq 1$ and $\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=12/n^{2}$. Therefore, for $n\geq 3$ the largest Eigen value of $\mathsf{D}\mathsf{D}^{\top}$ is at least $3$ whereas the smallest Eigen value is at most $\mathcal{O}(n^{-2})$. Thus, the condition number of the dual problem is $\Omega(n^{2})$ which renders dual ascent algorithms, and in particular, DPG slow to converge.
Appendix C Upper Bounds on Slack Variables
Denote the optimum of the original problem by ${\mathbf{x}}_{1:n}^{*}$, we claim that $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\leq\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$ and thus we can set $D_{i}=\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$. If $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}=0$, the bound clearly holds. Otherwise let $S=\{j:|x_{i+1,j}^{*}-x_{i,j}^{*}|=\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\}$. Consider $j\in S$ and let $s_{i,j}=\text{sign}(x_{i+1,j}^{*}-x_{i,j}^{*})$. From the optimality condition, there exists $\boldsymbol{\alpha}_{1:n-1}$ such that
$x_{i+1,j}^{*}-x_{i,j}^{*}$ $=$ $(y_{i+1,j}+\alpha_{i+1,j}-\alpha_{i,j})-(y_{i,j}+\alpha_{i,j}-\alpha_{i-1,j})$
$=$ $y_{i+1,j}-y_{i,j}+\alpha_{i+1,j}-2\alpha_{i,j}+\alpha_{i-1,j}~{}.$
In addition, we have $s_{i,j}\alpha_{i,j}\geq 0$. Multiplying both side by $s_{i,j}$ we get
$|x_{i+1,j}^{*}-x_{i,j}^{*}|$ $=$ $s_{i,j}(y_{i+1,j}-y_{i,j})+s_{i,j}\alpha_{i+1,j}+s_{i,j}\alpha_{i-1,j}-2|\alpha_{i,j}|$
$\leq$ $|y_{i+1,j}-y_{i,j}|+|\alpha_{i+1,j}|+|\alpha_{i-1,j}|-2|\alpha_{i,j}|~{}.$
Summing over $j\in S$ and dividing by $|S|$ we get,
$\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\;\leq\;\frac{1}{|S|}\big{(}\sum_{j\in S}|y_{i+1,j}-y_{i,j}|+\|\boldsymbol{\alpha}_{i+1}\|_{1}+\|\boldsymbol{\alpha}_{i-1}\|_{1}-2\big{)}\;\leq\;\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}~{},$
using the fact that for $j\notin S:\alpha_{i,j}=0$ and $\|\boldsymbol{\alpha}_{i}\|_{1}=1$ since $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\neq 0$.
Appendix D Iterative Algorithm for High-Order Variational Penalties
Note that the following modified smooth problem can be solved in linear time,
$\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\lambda_{i}|(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})|^{2}~{}~{}.$
Since the problem is differentiable everywhere, the optimality condition amounts to a linear system $\mathsf{D}{\mathbf{x}}=\mathbf{y}$ where $\mathsf{D}$ is a five-diagonal matrix. The system can be solved using Gauss elimination to reduce the matrix to an upper triangular one in linear time. Once triangulated, ${\mathbf{x}}$ is simply read out from the resulting system. Then, similar to Algorithm 1, we iteratively replace the non-smooth objective by its quadratic upper bound. Namely, on iteration $t$, we obtain an exact solution from,
${\mathbf{x}}^{t+1}=\displaystyle\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\frac{|(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})|^{2}}{2\rho_{i}^{t}}$
where $\rho_{i}^{t}=\displaystyle\max\big{(}|(x_{i+2}^{t}-x_{i+1}^{t})-(x_{i+1}^{t}-x_{i}^{t})|,\,\epsilon\big{)}$.
The line of proof as above can be carried out except that now we have,
$|(x_{i+2}-2x_{i+1}+x_{i})-(x^{\prime}_{i+2}-2x^{\prime}_{i+1}+x^{\prime}_{i})|^{2}\leq 4\left(|x_{i+2}-x^{\prime}_{i+2}|^{2}+2|x_{i+1}-x^{\prime}_{i+1}|^{2}+|x_{i}-x^{\prime}_{i}|^{2}\right)~{}.$
Summing over $i$ we end up with a factor of 16 instead of 4. This procedure can be generalized to $k$’th order finite difference penalties. Concretely, let us define the matrix by $\tilde{\mathsf{D}}$ the discrete differentiation matrix with zero padding,
$\tilde{\mathsf{D}}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\\ 0&0&0&\dots&0&0\\ \end{bmatrix}\in\mathsf{R}^{n\,\times\,n}~{}.$
We also employ this matrix sans the last row in Appendix B. The penalized $k$’th order finite differences we care to solve is,
$\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\big{\|}\big{[}\mathsf{\tilde{D}}^{k}{\mathbf{x}}\big{]}_{1:n-k}\big{\|}_{1}~{}.$ (17)
In terms of run time, we “pay” a factor of $4^{k}$ for the number of iterations required to converge and each iteration takes $\mathcal{O}(k^{2}n)$ for solving the linear system. In comparison, when using the dual proximal gradient method, the complexity of each iteration grows as $\mathcal{O}(\gamma^{k})$ where $\gamma=\Omega(n^{2})$ is the condition number of the first order problem. Pseudocode for $k>0$ is provided below.
Algorithm 2 $k^{\text{th}}$-Order Nariat
  initialize ${\mathbf{x}}^{0}=\mathbf{y}$
  for $t=1\ \textbf{to}\ T$ do
     $\mathsf{H}^{t}=\begin{bmatrix}\operatorname{diag}\bigg{(}\bigg{[}\Big{|}\big{[}\mathsf{\tilde{D}}^{k}\,{\mathbf{x}}^{t}\big{]}_{1:n-k}\Big{|}\bigg{]}_{\epsilon}\bigg{)}^{-1/2}&\mathsf{O}_{n-k\times k}\\ \mathsf{O}_{k\times n-k}&\mathsf{O}_{k\times k}\\ \end{bmatrix}$
     ${\mathbf{x}}^{t+1}=\displaystyle\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\frac{1}{2}\|\mathsf{H}^{t}\,\mathsf{\tilde{D}}^{k}\,{\mathbf{x}}\|^{2}$
  end for
  return $\mathbf{x}$
Figure 8: Python code of algorithms from Sec. 2.