### v2

# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

Abstract
We describe an apparatus for subgradient-following of the optimum of convex problems with variational penalties. In this setting, we receive a sequence $y_{i},\ldots,y_{n}$ and seek a smooth sequence $x_{1},\ldots,x_{n}$. The smooth sequence needs to attain the minimum Bregman divergence to an input sequence with additive variational penalties in the general form of $\sum_{i}{}g_{i}(x_{i+1}-x_{i})$. We derive known algorithms such as the fused lasso and isotonic regression as special cases of our approach. Our approach also facilitates new variational penalties such as non-smooth barrier functions.
We then introduce and analyze new multivariate problems in which ${\mathbf{x}}_{i},\mathbf{y}_{i}\in\mathbb{R}^{d}$ with variational penalties that depend on $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. The norms we consider are $\ell_{2}$ and $\ell_{\infty}$ which promote group sparsity. We also derive a novel lattice-based procedure for subgradient following of variational penalties characterized through the output of arbitrary convolutional filters. This paradigm yields efficient solvers for high-order filtering problems of temporal sequences in which sparse discrete derivatives such as acceleration and jerk are desirable.
1 Sub-Introduction
To start, let us examine two seemingly different problems, the fused lasso [25, 10] and isotonic regression [2, 4]. The input to the two procedures is a sequence $\mathbf{y}=y_{1},\ldots,y_{n}$ where the goal is two obtain smoothed sequences, denoted by ${\mathbf{x}}$, that are the solutions of the following optimization problems,
[…]
We can unify the two problems into a single abstract problem of the form,
$\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}g_{i}(x_{i}-x_{i+1})~{}~{}.$ (1)
Here, both $h_{i}:\mathbb{R}\to\mathbb{R}_{+}$ and $g_{i}:\mathbb{R}\to\mathbb{R}_{+}$ are convex in their single argument. Specific choices for $h_{i}$ and $g_{i}$ yield as special cases the fused-lasso, by setting $g_{i}(x_{i}-x_{i+1})=\lambda_{i}|x_{i}-x_{i+1}|$, and isotonic regression, by defining $g_{i}(x_{i}-x_{i+1})=0$ if $x_{i+1}\geq x_{i}$ and $\infty$ otherwise. Albeit the superficially different form of the variational penalty $g_{i}$, our apparatus facilitates gradient following procedures which differ by a single line of code as shown in Fig. 8. As in the case of isotonic regression, this procedure can facilitate non-smooth variational penalties for $g_{i}$. We fondly refer to the general subgradient-following paradigm as the Untangled Lariat. To illustrate the power of the Untangled Lariat we derive a novel derivative of the generic problem given by Eq. (1) which uses fairly general barrier functions for $g_{i}(\cdot)$.
We next generalize the Untangled Lariat to multivariate settings in which each element of the input sequence is a vector. We thus seek a solution ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$ where ${\mathbf{x}}_{i}\in\mathbb{R}^{d}$ for all $i\in[n]$. In order to obtain variational sparsity in the multivariate case we replace the absolute value $|x_{i+1}-x_{i}|$ with norm penalties over the differences $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. To start, we derive a closed-form for the squared $2$-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$. We use the closed form solution to build surrogate functions for the $2$-norm itself, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$, and the infinity-norm, $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$. The result is an iterative yet efficient solver for the problem,
$\min_{{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}}\>\sum_{i=1}^{n}h_{i}({\mathbf{x}}_{i})+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i+1}\|_{p}~{}\mbox{ where }~{}p\in\{2,\infty\}.$
For brevity of the derivation of the multivariate setting we confine ourselves to the case where $h_{i}({\mathbf{x}}_{i})=\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}$.
Last, but not least, we introduce a construction from signal processing into our setting by penalizing the convolution of ${\mathbf{x}}$ with a predefined convolution filter $\boldsymbol{\tau}$. Concretely, we provide an efficient procedure for finding the optimum of,
$\min_{\mathbf{x}}\sum_{i}h_{i}(x_{i})+\sum_{i}\big{|}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{|}~{}~{},$
where $\boldsymbol{\tau}\in\mathbb{R}^{k}$ is a predefined yet arbitrary filter of $k-1$ free parameters. We introduce a subgradient-following procedure which operates over a $k$-dimensional lattice. It takes, in the worst case, $\mathcal{O}(n^{k})$ operations to construct the lattice from which the optimal solution is obtained. In applications where $\mathbf{y}$ represents a temporal sequence, the variational penalty of the fused lasso $|x_{i+1}-x_{i}|$ promotes smoothed solutions with regions of zero discrete “velocity”. Analogously, the high-order Untangled Lariat is a filtering and smoothing apparatus for problems in which we seek regions of zero acceleration, zero jerk, and other forms of finite differences.
The initial motivation for this work can be traced back to the influential work on the fused lasso [25, 10] and isotonic regression [2, 4, 29]. The setting of the fused lasso has been revisited numerous times and found various venues for applications. See for instance [12, 13], though any attempt to be exhaustive here is going to do injustice to the plethora of applications that employ the fused lasso. Several authors [8, 16, 18] described and analyzed a linear time algorithm for solving the fused lasso problem which is morally equivalent to Untangled Lariat with the aforementioned specific choices for $h_{i}$ and $g_{i}$. Alternative algorithms for the fused lasso and similar problems, such as the ones described in [3, 7], do not guarantee a runtime linear in $n$ nonetheless demonstrate excellent empirical results.
The paper that is morally most relevant to our work is [18] where the authors follow the derivation in [16] to generalize the algorithm to non-smooth unary functions and to trees. The procedure described in [18] is an instance of a primal-dual message passing algorithm [11]. The authors also consider unary functions subject to a pairwise penalty term which also include the fused lasso penalty as a special case. The lens of the work presented in [18] is probabilistic in nature and does not consider general pairwise penalties.
Several other forms of variational penalties were the focus specific extension of the fused lasso and isotonic regression. For instance, the works presented in [29] considers isotonic regression with $\ell_{1}$ norm for the variational penalty. We readily obtain the same procedure as a special case of our approach. The problem of high-order variational penalties was originally proposed by [17] and its statistical properties were studied by [24, 27]. An algorithmic perspective was provided in [17] which discusses an approximate solution using a primal-dual interior point method. For penalties of the form $\lambda\sum_{i}|x_{i+1}-x_{i}|$, the entire solution path for admissible values of $\lambda$, is provided using a dual path algorithm in [26, 1]. To the best of our knowledge, we give the first efficient analytic solution for a given $\lambda$.
We generalize the aforementioned research papers by providing a single unified algorithmic paradigm for general convex unary functions and a broad class of pairwise terms. Our analysis technique deviates from prior research by employing the Fenchel dual of the primal problem, a view that was advocated in the context of online learning by Shalev-Shwartz [22]. For a thorough study of Fenchel duality and its connection to Lagrange dual see for instance Chapter 3 in [BorweinLewis06]. The incorporation of Fenchel duality leads to a simple and useful generalization to Bregman divergences [6] between ${\mathbf{x}}$ and $\mathbf{y}$ as the objective. This generalization does not bear an additional computation cost. We provide in an appendix Python code that implements subgradient following procedures for most of the settings discussed in this paper.
2 Abstract Subgradient Following Algorithm
Let us now consider a general minimization with sequential penalties of the form,
$U=\min_{{\mathbf{x}}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}g_{i}(x_{i}-x_{i+1})~{}~{}.$ (2)
Throughout the paper we assume that $h_{i}$ and $g_{i}$ are convex and lower semi-continuous. The Fenchel dual of a function $f$ is defined as,
$f^{*}(\alpha)=\sup_{x\in\mathrm{dom}(f)}\big{\{}\alpha x-f(x)\big{\}}~{}~{}.$
The domain of the functions we consider is closed, compact, and convex. We can thus replace the supremum above with a maximum. We next utilize Fenchel-Moreau theorem for each $g_{i}$,
$g_{i}(x_{i}-x_{i-1})=\max_{\alpha}\big{\{}\alpha(x_{i}-x_{i-1})-g_{i}^{*}(\alpha)\big{\}}~{}~{},$
in order to evaluate $U$ as follows,
$U$ $=$ $\min_{{\mathbf{x}}}\,\max_{\boldsymbol{\alpha}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]~{}~{},$
We confine ourselves to settings for which strong duality holds thus,
$U$ $=$ $\min_{x_{1}}\,\max_{\alpha_{1}}\,\min_{x_{2}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=1}^{n-1}\left[\alpha_{i}(x_{i}-x_{i+1})-g_{i}^{*}(\alpha_{i})\right]~{}~{}.$
To find the optimum we define a sequence of partial single variable functions defined as,
$f_{i}(x_{i})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(x_{j}-x_{j+1})-g_{j}^{*}(\alpha_{j})\right]~{}~{}.$ (3)
For boundary conditions we let $\forall x:f_{n}(x)=0$. In addition, we define an auxiliary function $f_{i-\frac{1}{2}}(x_{i})=f_{i}(x_{i})+h_{i}(x_{i})$. Using these definitions we have $U=\min_{x_{1}}f_{\frac{1}{2}}(x_{1})$. Next we introduce a summation-free form for $f_{i}$,
$f_{i}(x_{i})$ $=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+1}(x_{i+1})+h_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$ (4)
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+\frac{1}{2}}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}x_{i}-g_{i}^{*}(\alpha_{i})-f_{i+\frac{1}{2}}^{*}(\alpha_{i})~{}~{}.$
Note that to find $f_{i}(x_{i})$ we need to obtain the minimum of $f_{i+\frac{1}{2}}(x_{i+1})-\alpha_{i}x_{i+1}$ with respect to $x_{i+1}$. Using this abstraction, obtaining the optimum for $x_{i+1}$ “only” requires the knowledge of $\alpha_{i}$. Analogously, to calculate $f_{i}$ itself we need to find the optimum of $\alpha{}x_{i}-g_{i}^{*}(\alpha)-f_{i+\frac{1}{2}}^{*}(\alpha)$ with respect to $\alpha$ which gives us $\alpha_{i}$. This in turn only requires the knowledge of $x_{i}$. To recap, the optimization sub-problem for $x_{i+1}$ solely uses the information on $\alpha_{i}$ and analogously the optimization sub-problem for $\alpha_{i}$ only requires information of $x_{i}$. Let $\mathsf{x}_{1},\mathsf{a}_{1},\mathsf{x}_{2},\mathsf{a}_{2},\ldots,\mathsf{a}_{n-1},\mathsf{x}_{n}$ denote a sequence of locally-admissible pairs $\mathsf{x}_{i},\mathsf{a}_{i}$ and $\mathsf{a}_{i},\mathsf{x}_{i+1}$ where the full sequence is not necessarily globally optimal. The subgradient conditions for pair-restricted optimality define two admissible sets,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1})\big{\}}$ (5)
$S_{i}$ $=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})\big{\}}~{}~{}.$ (6)
Given $S_{i+1}=\left\{(\mathsf{x}_{i+1},\mathsf{a}_{i+1}):\mathsf{a}_{i+1}\in\partial f_{i+1}(\mathsf{x}_{i+1})\right\}$ we perform an expansion,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i+1}(\mathsf{x}_{i+1})+\partial h_{i+1}(\mathsf{x}_{i+1})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i+1},\mathsf{a}_{i}):\mathsf{a}_{i}-\mathsf{a}_{i+1}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i+1})\in S_{i+1}\big{\}}~{}~{}.$
We next derive a recursive map $S_{i+1}\mapsto S_{i}$ by going through $S_{i+\frac{1}{2}}$. From the representation of $f_{i}(x_{i})$ as given by Eq. (4) we get that the pair $(\mathsf{x}_{i},\mathsf{a}_{i})$ is admissible when
$\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial f_{i+\frac{1}{2}}^{*}(\mathsf{a}_{i})~{}~{}.$
Starting with the definition of $S_{i}$ and expanding it based on the above property we get,
$S_{i}$ $=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{a}_{i}\in\partial f_{i}(\mathsf{x}_{i})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}\in\partial g_{i}^{*}(\mathsf{a}_{i})+\partial{f_{i+\frac{1}{2}}}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\mathsf{x}_{i},\mathsf{a}_{i}):\mathsf{x}_{i}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}~{}.$
For the last equality we used the fact that,
$(\mathsf{x}_{i+1},\mathsf{a}_{i})\in S_{i+\frac{1}{2}}~{}\Leftrightarrow~{}\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial f^{*}_{i+\frac{1}{2}}(\mathsf{a}_{i})~{}~{}.$
In summary, we obtain the following aesthetically succinct recursion,
$\begin{array}[]{lcl}S_{i+\frac{1}{2}}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{a}-\mathsf{a}^{\prime}\in\partial h_{i+1}(\mathsf{x})\;;\;(\mathsf{x},\mathsf{a}^{\prime})\in S_{i+1}\right\}\\ S_{i}&=&\left\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;(\mathsf{x}^{\prime},\mathsf{a})\in S_{i+\frac{1}{2}}\right\}~{}~{}.\end{array}$
To kick-start the recursive procedure we set $S_{n}$ as a boundary condition,
$S_{n}\equiv\big{\{}(\mathsf{x},0):\mathsf{x}\in\text{dom}(h_{n})\big{\}}~{}~{}.$
We end the recursion upon reaching $S_{\frac{1}{2}}$ and then set $x_{1}$ to be the left argument of the pair $(\mathsf{x},0)\in S_{\frac{1}{2}}$. Once $x_{1}$ is inferred we can read out the rest of the sequence, $x_{i}$ for $i>1$, by unravelling the recurrence,
[…]
3 Derived Algorithms
In this section we apply the subgradient following framework to specific settings. We first show that the fused lasso is a simple instance. We next derive other variational (fused) penalties, including barrier penalties that were not analyzed before. We also show that the squared distance between ${\mathbf{x}}$ and $\mathbf{y}$ can be generalized at no additional computational cost to separable Bregman divergences.
3.1 Revisiting the Fused Lasso
To recover the Fused Lasso algorithm using the subgradient-following method we choose, $h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}$ and $g_{i}(\delta)=\lambda_{i}|\delta|$. Using basic calculus these choices imply that $\nabla h_{i}(x)=x-y_{i}$ and $g_{i}^{*}(\alpha)=\mathbf{1}_{|\alpha|\leq\lambda_{i}}$. Over the domain of $g_{i}^{\star}$, namely $|\alpha|\leq\lambda_{i}$, its subgradient is,
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=-\lambda_{i}\\ 0&|\alpha|<\lambda_{i}\\ [0,\infty)&\alpha=\lambda_{i}\end{cases}~{}~{}.$
We next show that $\forall i\in[n]$, there exists a mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$, where $\mathsf{a}_{i}(\cdot)$ is a continuous non-decreasing piecewise linear function in $\mathsf{x}$. This property serves as a succinct representation of $S_{i}=\{(\mathsf{x},\mathsf{a}_{i}(\mathsf{x}))\}$ which takes the following recursive form,
$\mathsf{a}_{i}(\mathsf{x})=\Big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}~{}.$ (7)
To start, let us define $a_{n}(\mathsf{x})\equiv 0$ which adheres with the definition of $S_{n}$ from the previous section. We next show that the form of $S_{i}$ holds inductively, using the definitions of $S_{i}$ and $S_{i+\frac{1}{2}}$ from the previous section. Assume that Eq. (7) holds true for $S_{i+1}$. The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$S_{i+\frac{1}{2}}=\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):\mathsf{x}\in\mathsf{R}\}~{}~{}.$
For convenience we define $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. In order to evaluate,
$S_{i}=\{(\mathsf{x},\mathsf{a}):\mathsf{x}-\mathsf{x}^{\prime}\in\partial g_{i}^{*}(\mathsf{a})\;;\;\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})\}$
we need to analyze $\Delta\mathsf{x}=\mathsf{x}-\mathsf{x}^{\prime}$ for the admissible set of values of $\mathsf{a}$ under the map $g_{i}^{*}$ which is $[-\lambda_{i},\lambda_{i}]$. Since $\mathsf{a}_{i+1}(\cdot)$ is continuous and non-decreasing, it implies that $\mathsf{a}_{i}^{\prime}(.)$ is strictly-increasing and bijective. Therefore, there exists two unique boundary values $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=-\lambda_{i}$ and $z_{i}^{+}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{+})=+\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})$ such that $-\lambda_{i}<\mathsf{a}<\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=\{0\}$. Thus, $\mathsf{x}=\mathsf{x}^{\prime}$ and both are in $(z_{i}^{-},z_{i}^{+})$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\in(z_{i}^{-},z_{i}^{+})$ yields,
$S_{i}\cap\{\mathsf{x}:z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$ $=$ $\{(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x})):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}$
$=$ $\{(\mathsf{x},\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}):z_{i}^{-}<\mathsf{x}<z_{i}^{+}\}~{}~{}.$
We last need to address the cases $\mathsf{a}=\pm\lambda_{i}$. For $\mathsf{a}=\mathsf{a}_{i}^{\prime}(z_{i}^{+})=\lambda_{i}$ we have $\partial g^{*}_{i}(\mathsf{a})=[0,\infty)$. Thus, for any $\mathsf{x}\geq z_{i}^{+}$ we have $\mathsf{x}-z_{i}^{+}\in\partial g^{*}_{i}(\lambda_{i})$ and therefore $(\mathsf{x},\lambda_{i})\in S_{i}$. Analogously, for $\mathsf{a}=-\lambda_{i}$ which implies that for $\mathsf{x}\leq z_{i}^{-}$, it holds that $\mathsf{x}-z_{i}^{-}\in\partial g^{*}_{i}(-\lambda_{i})\Rightarrow(\mathsf{x},-\lambda_{i})\in S_{i}$. We therefore get that the restriction of $S_{i}$ to $\mathsf{x}\geq z_{i}^{+}$ and $\mathsf{x}\leq z_{i}^{-}$ yields,
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\geq z_{i}^{+}\}$ $=$ $\{(\mathsf{x},\lambda_{i}):\mathsf{x}\geq z_{i}^{+}\}$
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},-\lambda_{i}):\mathsf{x}\leq z_{i}^{-}\}~{}~{}.$
Since $\mathsf{a}_{i}^{\prime}(\cdot)$ is increasing, we have for $\mathsf{x}\geq z_{i}^{+}\Rightarrow\mathsf{a}_{i}^{\prime}(\mathsf{x})\geq\lambda_{i}$ and $\mathsf{x}\leq z_{i}^{-}\Rightarrow\mathsf{a}_{i}^{\prime}(\mathsf{x})\leq-\lambda_{i}$. Therefore, we established that
$S_{i}=\left\{\Big{(}\mathsf{x},\Big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}\Big{)}:\mathsf{x}\in\mathsf{R}\right\}~{}~{}.$
Figure 1: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for the fused lasso.
3.2 Sparse Fused Lasso
We next combine both the standard lasso and fusion penalties. Namely, we choose
$h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}+\beta_{i}|x|~{}\mbox{ and }~{}g_{i}(\delta)=\lambda_{i}|\delta|~{}~{}.$
These choices imply that
$\partial h_{i}(x)=\begin{cases}x-y_{i}+\beta_{i}&x>0\\ [-y_{i}-\beta_{i},-y_{i}+\beta_{i}]&x=0\\ x-y_{i}-\beta_{i}&x<0\end{cases}~{}~{}.$
Generalizing the fused lasso, the succinct representation for the sparse fused lasso is of the form,
$S_{i}=\{(\mathsf{x},\mathsf{a}_{i}^{+}(\mathsf{x})):\mathsf{x}>0\}\,\cup\,\{(0,\mathsf{a}):\mathsf{a}_{i}^{-}(0)\leq\mathsf{a}\leq\mathsf{a}_{i}^{+}(0)\}\,\cup\,\{(\mathsf{x},\mathsf{a}_{i}^{-}(\mathsf{x})):\mathsf{x}<0\}~{}~{}.$
The locally-admissible set consists of three disjoint subsets induced by the sign of the first argument $\mathsf{x}$. It is immediate to verify that the subgradient at any value but zero is unique. Hence, the mappings for positive and negative numbers are well-defined and take similar recursive forms to that of the fused lasso,
$\mathsf{a}_{i}^{\pm}(\mathsf{x})=\Big{[}\mathsf{a}_{i+1}^{\pm}(\mathsf{x})+\mathsf{x}-y_{i+1}\pm\beta_{i+1}\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}~{}.$
The sole value that requires further attention is at $\mathsf{x}=0$. The admissible interval at zero is determined by the two boundary values, $\mathsf{a}_{i}^{-}(0)$ and $\mathsf{a}_{i}^{+}(0)$. To facilitate an efficient implementation, we store the value of $\Delta_{i}=\mathsf{a}_{i}^{+}(0)-\mathsf{a}_{i}^{-}(0)$. Prior to rectification, we update $\Delta_{i}\leftarrow\Delta_{i-1}+2\beta_{i}$. Analogous to the fused lasso, we traverse the segments of $\mathsf{a}_{i}(\cdot)$ left-to-right until we reach the segment encapsulating $-\lambda_{i}$. The only difference is that we increase the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero. Analogously, we decrease the value of $\mathsf{a}_{i}(\cdot)$ by $\Delta_{i}$ upon crossing the zero when traversing right-to-left. After obtaining the critical points $z_{i}^{\pm}$, we update $\Delta_{i}$ accordingly. In each iteration, we incur constant additional costs and thus the total runtime remains intact as $\mathcal{O}(n)$.
Figure 2: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for the sparse fused lasso.
3.3 Bregman Divergences
Before proceeding with nascent total variation regularizers, we now describe a simple way for incorporating convex losses for $h_{i}(\cdot)$ at zero additional computational cost. To remind the reader, given a strictly convex function $\phi:\mathbb{R}\to\mathbb{R}$, the Bregman divergence between $x\in\mathbb{R}$ and $y\in\mathbb{R}$ is defined as, $D_{\phi}(x,y)=\phi(x)-(\phi(y)+\nabla\phi(y)(x-y))$. Our generalization amount to setting, $h_{i}(x)=D_{\phi}(x,y_{i})$ and keeping $g_{i}(\delta)$ intact, thus retaining for the fused lasso $g_{i}(\delta)=\lambda_{i}|\delta|$. This choice for $h_{i}$ implies that $\nabla h_{i}(x)=\nabla\phi(x)-\nabla\phi(y_{i})$. Since $\phi(\cdot)$ is strictly convex, $\nabla\phi(\cdot)$ is continuous and strictly increasing. Following the same line of derivation as above we get that the recursive mappings for each $\mathsf{a}_{i}(\cdot)$ takes the following form,
$\mathsf{a}_{i}(\mathsf{x})=\Big{[}\mathsf{a}_{i+1}(\mathsf{x})+\nabla\phi(\mathsf{x})-\nabla\phi(y_{i+1})\Big{]}_{-\lambda_{i}}^{+\lambda_{i}}~{}~{}.$ (8)
This implies that $\mathsf{a}_{i}(.)$ is a piecewise linear function of $\nabla\phi(.)$. We thus follow the same procedure above whilst replacing $\mathsf{x}\mapsto\nabla\phi(x)$ for each linear segment in $\mathsf{a}_{i}$. This generalization using a Bregman divergence is applicable throughout the remainder of the section. We keep using the squared error for brevity of our derivations.
3.4 Isotonic Regression
In isotonic regression the goal is to find the best non-decreasing sequence, $x_{1}\leq x_{2}\leq\ldots\leq x_{i}\leq x_{i+1}\leq\ldots\leq x_{n}$ for which the sum of the losses $h_{i}(x)=\frac{1}{2}(x-y_{i})^{2}$ is minimized. We replace the fused penalties with $g_{i}(\delta)=\mathbf{1}_{\delta\leq 0}$. This choice implies that $g_{i}^{*}(\alpha)=\mathbf{1}_{\alpha\geq 0}$ and its subgradient is,
$\partial g_{i}^{*}(\alpha)=\begin{cases}(-\infty,0]&\alpha=0\\ 0&\alpha>0\end{cases}~{}~{}.$
Parallel to the fused lasso, there exists a succinct representation for isotonic regression of the form $S_{i}=\{(\mathsf{x},\mathsf{a}_{i}(\mathsf{x})):\mathsf{x}\in\mathsf{R}\}$ which is prescribed recursively as,
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{+}$
where $[z]_{+}=\max(z,0)$. The derivation follows the same lines as above as there exists a unique boundary value $z_{i}^{-}\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z_{i}^{-})=0$. The resulting description of $S_{i}$ is obtained from the union of,
$S_{i}\cap\{\mathsf{x}:\mathsf{x}>z_{i}^{-}\}$ $=$ $\{(\mathsf{x},\alpha_{i}^{\prime}(\mathsf{x})):\mathsf{x}>z_{i}^{-}\}$
$S_{i}\cap\{\mathsf{x}:\mathsf{x}\leq z_{i}^{-}\}$ $=$ $\{(\mathsf{x},0):\mathsf{x}\leq z_{i}^{-}\}~{}~{}.$
We can in fact entertain a simple generalization which encompasses both absolute difference penalties, isotonic regression, and asymmetric difference penalties by defining,
$g_{i}(\delta)=\bar{\lambda}_{i}[\delta]_{+}\,-\,\underline{\lambda}_{i}[-\delta]_{+}\,\mbox{ with }\,\bar{\lambda}_{i}\geq 0\,,\,\underline{\lambda}_{i}\leq 0~{}~{}.$
The recursive mapping takes the following form,
$\mathsf{a}_{i}(\mathsf{x})=\big{[}\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}\big{]}_{\underline{\lambda}_{i}}^{\bar{\lambda}_{i}}~{}~{}.$
Setting $-\underline{\lambda}_{i}=\bar{\lambda}_{i}=\lambda_{i}$ yields the fused lasso whereas choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\infty$ provides a template for isotonic regression. In addition, choosing $\underline{\lambda}_{i}=0$ and $\bar{\lambda}_{i}=\lambda$ recovers the setting of [28]. To conclude this section we would like to note that isotonic regression can be used as the core for finding the closest vector to $\mathbf{y}$ subject to a $\leq k$-modality constraint. Due to submodularity of the objective the total runtime would be $\mathcal{O}(kn)$.
Figure 3: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for isotonic regression.
3.5 Fused Barriers
For certain data analysis applications the goal is to eliminate outlier points in an observed sequence rather than smooth or filter the sequence. We cast the problem as requiring the differences between consecutive points not to exceed prescribed maximal changes $b_{i}$ by establishing a series of barrier functions, $g_{i}(\delta)=\mathbf{1}_{|\delta|\leq\lambda_{i}}$. Each $b_{i}$ here can be thought of as the maximum tolerance to changes and is not necessarily small and thus restrictive. For this choice of $g_{i}$ we get,
$\partial g_{i}^{*}(\alpha)=\begin{cases}-\lambda_{i}&\alpha<0\\ [-\lambda_{i},\lambda_{i}]&\alpha=0\\ \lambda_{i}&\alpha>0\end{cases}~{}~{}.$ (9)
Albeit being slightly more involved, we next show that there exists a succinct representation of $S_{i}$ in the form of a piecewise linear map, $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. First, let us reuse the intermediate function $\mathsf{a}_{i}^{\prime}$ from the derivation of the fused lasso algorithm, $\mathsf{a}_{i}^{\prime}(\mathsf{x})=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$. We denote by $\kappa_{\lambda}(\cdot)$ the Kvetsh operator (from Yiddish, literally “to squeeze, shrink”), colloquially called shrinkage,
$\kappa_{\lambda}(z)=\text{sign}(z)\,\big{[}|z|-\lambda\big{]}_{+}~{}~{}.$
The construction of $\mathsf{a}_{i}$ from $\mathsf{a}_{i+1}$ consists of the following steps,
$\mathsf{a}_{i}^{\prime}(\mathsf{x})$ $:=\mathsf{a}_{i+1}(\mathsf{x})+\mathsf{x}-y_{i+1}$ (Add Coordinate)
$z_{i}$ $:=z\,\;\text{s.t.}\;\mathsf{a}_{i}^{\prime}(z)=0$ (Zero Crossing)
$\mathsf{a}_{i}(\mathsf{x})$ $:=\mathsf{a}_{i}^{\prime}\big{(}z_{i}+\kappa_{\lambda_{i}}(\mathsf{x}-z_{i})\big{)}$ (Kvetsh)
Informally, after constructing the intermediate function $\mathsf{a}^{\prime}_{i}$, which is strictly monotone, we find the point $z_{i}$ where $\mathsf{a}^{\prime}_{i}$ crosses zero. We then insert an interval of zero slope whose center is $z_{i}$ and its support spans from $z_{i}-\lambda_{i}$ through $z_{i}+\lambda_{i}$.
We now provide the details of the derivation of $\mathsf{a}_{i}$. Since $z_{i}$ is the locus of the zero of $\mathsf{a}^{\prime}_{i}$ then for $\mathsf{x}^{\prime}>z_{i}$ we have $\mathsf{a}=\mathsf{a}_{i}^{\prime}(\mathsf{x}^{\prime})>0\Rightarrow\partial g_{i}^{*}(\mathsf{a})=\lambda_{i}$ based on Eq. (9). Thus, for any $\mathsf{x}=\mathsf{x}^{\prime}+\lambda_{i}>z_{i}+\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\in S_{i}$. Analogously, for $\mathsf{x}<z_{i}-\lambda_{i}$ we have $(\mathsf{x},\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\in S_{i}$. Last, for $(z_{i},\mathsf{a}^{\prime}_{i}(z_{i}))=(z_{i},0)$ we have $\partial g_{i}^{*}(0)=[-\lambda_{i},\lambda_{i}]$. Therefore, for any $\mathsf{x}$ such that $z_{i}-\lambda_{i}\leq\mathsf{x}\leq z_{i}+\lambda_{i}$ it holds that $\mathsf{x}-z_{i}\in\partial g_{i}^{*}(0)\Rightarrow(\mathsf{x},0)\in S_{i}$. The resulting $S_{i}$ is obtained from the union of,
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}-\lambda_{i}))\,:\,\mathsf{x}<z_{i}-\lambda_{i}\}$
[…]
$S_{i}\cap\{\mathsf{x}\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}$ $=$ $\{(\mathsf{x},\,\mathsf{a}_{i}^{\prime}(\mathsf{x}+\lambda_{i}))\,:\,\mathsf{x}>z_{i}+\lambda_{i}\}~{}~{}.$
Eq. (Kvetsh) provides a succinct functional form of the union of these three sets.
Since each $\mathsf{a}_{i}(\cdot)$ is a piecewise linear function we can employ the same algorithmic skeleton used for the fused lasso. Alas, in order to construct $\mathsf{a}_{i}$ we need to traverse $\mathsf{a}^{\prime}_{i}$ and find the locus, $z_{i}$, of its zero. This search would require $\mathcal{O}(n)$ time and thus the total run time would amount to $\mathcal{O}(n^{2})$. We provide in Appendix A a description of a more efficient implementation which uses a more elaborated data structure in order to reduce the amortized runtime to $\mathcal{O}(n\log(n))$.
Figure 4: Construction of $\mathsf{a}_{3}(\cdot)$ from $\mathsf{a}_{4}(\cdot)$ and optimal solution for fused barriers.
We conclude the section with a short discussion and illustration of the characteristics of the variational penalties of the fused lasso by contrasting it with barrier constraints on the variation. To this end, we generated a sequence based on a quadratic function, $y_{i}=(i-50)^{2}/100$. We then contaminated the sequence with noise. We first added i.i.d noise to each element sampled from the uniform distribution over $[-\epsilon,\epsilon]$. We next added “shock” noise by replacing the value of a few elements chosen at random with a small negative number. The generating and noisy sequences are plotted on the left hand-side of Figure 5. We then solved least squares approximation for a range of variational penalties and constraints by varying $\lambda$. For each solution ${\mathbf{x}}(\lambda)$ we calculated the $\ell_{1}$ and $\ell_{\infty}$ variation norm. The results are shown on the right hand-side of Figure 5. The dashed two-sided arrows designate solutions attaining the same approximation error by the two methods. It is evident that for the same error the fused lasso naturally obtains a lower $\ell_{1}$ variation norm and it is more resilient to the uniform noise. In contrast, posing a barrier constraint on the variation provides a uniformly lower $\ell_{\infty}$ norm of the variation and is more resilient to shock noise.
Figure 5: Comparison least-squares sequence approximation with variational penalty of the fused lasso (FL) versus barrier constraints (FB).
4 Variationally Penalized Multivariate Regression
We next extend the subgradient following framework to multivariate settings. For concreteness we examine objectives for instance $i$ of the form,
$h_{i}({\mathbf{x}})\,=\,\frac{1}{2}\big{\|}{\mathbf{x}}-\mathbf{y}_{i}\big{\|}^{2}_{{Q}_{i}}\,=\,\frac{1}{2}({\mathbf{x}}-\mathbf{y}_{i})^{\top}\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}~{}{\mathbf{x}},\mathbf{y}_{i}\in\mathbb{R}^{n}~{}~{}.$ (10)
We assume w.l.o.g that the quadratic form is strictly positive definite, $\mathsf{Q}_{i}\succ 0$. The multivariate setting is intrinsically more complex than the scalar case due to the penalty entanglement of ${\mathbf{x}}_{i}$ with itself and with ${\mathbf{x}}_{i+1}$. To mitigate this difficulty we first examine the squared Euclidean distance between ${\mathbf{x}}_{i}$ and ${\mathbf{x}}_{i+1}$,
$g_{i}(\boldsymbol{\delta})\,=\,\frac{1}{2}\lVert{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\rVert^{2}\,=\,\frac{1}{2}\lVert\boldsymbol{\delta}\rVert^{2}~{}~{}.$
This setting generalizes the squared fused penalty of the scalar case for which $Q_{i}=[q_{i}\in\mathbb{R}]$ and thus can be absorbed into $\lambda_{i}$. With these assumptions we get,
$\nabla h_{i}({\mathbf{x}})=\mathsf{Q}_{i}({\mathbf{x}}-\mathbf{y}_{i})~{}\mbox{and}~{}\nabla g_{i}^{*}(\boldsymbol{\delta})=\boldsymbol{\delta}~{}~{}.$
The derivation of the subgradient following for this setting would also serve us as a building block in more complex multivariate problems described and analyzed in the sequel.
The Fenchel dual for the multivariate case where $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined as,
$f^{*}(\boldsymbol{\alpha})=\sup_{{\mathbf{x}}\in\mathrm{dom}(f)}\big{\{}\boldsymbol{\alpha}^{\top}{\mathbf{x}}-f({\mathbf{x}})\big{\}}~{}~{}.$
Both $h_{i}(\cdot)$ and $g_{i}^{*}(\cdot)$ defined above are strictly convex, as we require $\mathsf{Q}_{i}\succ 0$, thus the Fenchel duals for all $i\in[n]$ have a unique maximizer that forms dual feasible pairs $({\mathbf{x}},\boldsymbol{\alpha})$. The equation $\boldsymbol{\alpha}=\nabla\!f({\mathbf{x}})$ is typically referred to as the link function or mirror map. We use this property in the ensuing derivation.
Since the components of the gradients are of unary form and linear, the subgradient sets can be written succinctly as,
$S_{i}=\left\{({\mathbf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}{\mathbf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+{\mathbf{c}}_{i}=\mathbf{0}\right\}~{}~{}.$
Furthermore, as the mirror map here is bijective, the admissible sets of Eq. (5) and Eq. (6) can be written as,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!f_{i+\frac{1}{2}}(\boldsymbol{\mathsf{x}}_{i+1})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):\boldsymbol{\mathsf{a}}_{i}=\nabla\!f_{i}(\boldsymbol{\mathsf{x}}_{i})\big{\}}~{}~{}.$
We therefore effortlessly unravel the recursive form for $\mathsf{A}_{i-1}$, $\mathsf{B}_{i-1}$, ${\mathbf{c}}_{i-1}$ from $\mathsf{A}_{i}$, $\mathsf{B}_{i}$, ${\mathbf{c}}_{i}$,
$S_{i-\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{a}}-\boldsymbol{\mathsf{a}}^{\prime}=\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i});\;(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}^{\prime})\in S_{i}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\mathsf{A}_{i}\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\left(\boldsymbol{\mathsf{a}}-\mathsf{Q}_{i}(\boldsymbol{\mathsf{x}}-\mathbf{y}_{i})\right)+{\mathbf{c}}_{i}=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{}~{},$ $\vskip 3.0pt plus 1.0pt minus 1.0ptS_{i-1}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{x}}^{\prime}=\boldsymbol{\mathsf{a}};\;(\boldsymbol{\mathsf{x}}^{\prime},\boldsymbol{\mathsf{a}})\in S_{i-\frac{1}{2}}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})(\boldsymbol{\mathsf{x}}-\boldsymbol{\mathsf{a}})+\mathsf{B}_{i}\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}},\boldsymbol{\mathsf{a}}):(\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i})\boldsymbol{\mathsf{x}}+(\mathsf{B}_{i}(\mathsf{I}+\mathsf{Q}_{i})-\mathsf{A}_{i})\boldsymbol{\mathsf{a}}+({\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i})=\mathbf{0}\big{\}}~{}~{}.$
In summary, we get that,
$\mathsf{A}_{i-1}$ $=$ $\mathsf{A}_{i}-\mathsf{B}_{i}\mathsf{Q}_{i}$
$\mathsf{B}_{i-1}$ $=$ $\mathsf{B}_{i}(\mathsf{I}+\mathsf{Q}_{i})-\mathsf{A}_{i}$
${\mathbf{c}}_{i-1}$ $=$ ${\mathbf{c}}_{i}+\mathsf{B}_{i}\mathsf{Q}_{i}\mathbf{y}_{i}~{}~{}.$
For boundary conditions we require,
$(\mathsf{A}_{n},\mathsf{B}_{n},{\mathbf{c}}_{n})=(\mathsf{O},\,\mathsf{I},\,\mathbf{0})~{}\mbox{ and }~{}(\mathsf{A}_{1}-\mathsf{B}_{1}\mathsf{Q}_{1}){\mathbf{x}}_{1}+({\mathbf{c}}_{1}+\mathsf{B}_{1}\mathsf{Q}_{1}\mathbf{y}_{1})=0~{}~{}.$
Upon obtaining ${\mathbf{x}}_{1}$ along with $\boldsymbol{\alpha}_{0}\equiv\mathbf{0}$, we construct the trajectory of the solution using the backward recursion,
$\boldsymbol{\alpha}_{i}$ $=$ $\boldsymbol{\alpha}_{i-1}-\mathsf{Q}_{i}({\mathbf{x}}_{i}-\mathbf{y}_{i})$
${\mathbf{x}}_{i+1}$ $=$ ${\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}~{}~{}.$
5 Sparse Variationally Penalized Multivariate Regression
We henceforth use ${\mathbf{x}}_{1:n}$ to denote the sequence of vectors ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}$. While the algorithm of Sec. 4 entertains a concise form, it does not yield variational sparsity. To do so we can readily use the $1$-norm as the variational penalty. This would yield though element-wise fusion whereas full variational sparsity entails that for some indices $i\in[n]$, ${\mathbf{x}}_{i}={\mathbf{x}}_{i+1}$. In this section we use the algorithm from the previous section as a building block while replacing the norm-squared penalty $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}$ with either the $2$-norm $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$ of the difference vector or its infinity norm $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}$. For brevity, we assume throughout this section that $\forall i,\lambda_{i}=1$ and are thus omitted.
5.1 Multivariate Regression with $\ell_{2}$-Variational Penalty
When we replace each of the norm-squared variational penalties with the $2$-norm, the goal is to find the minimizer of,
$\phi({\mathbf{x}}_{1:n})=\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|~{}~{}.$ (11)
Alas, this formulation does not entail a closed form subgradient following procedure. The approach that we take here is to construct a relaxation that yields the following convergent algorithm.
Algorithm 1 Surrogate for Multivariate Untangled Lariat
  initialize ${\mathbf{x}}_{1:n}^{0}=\mathbf{y}_{1:n}$
  for $t=1\ \textbf{to}\ t_{\text{max}}$ do
     $\forall i:\rho_{i}^{t}=\max\big{(}\|{\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t}\|,\,\epsilon\big{)}$
     ${\mathbf{x}}^{t+1}_{1:n}=\displaystyle\operatorname*{\arg\!\min}_{{\mathbf{x}}_{1:n}}\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\frac{\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|^{2}}{2\rho_{i}^{t}}$
  end for
  return ${\mathbf{x}}^{t_{\text{max}}}_{1:n}$
We next show that Algorithm 1 converges linearly to the minimum of,
$\hat{\phi}({\mathbf{x}}_{1:n})=\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})~{}~{}.$ (12)
where $g_{\epsilon}(\cdot)$ is the Huber [15] loss defined as,
$g_{\epsilon}(\boldsymbol{\delta})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}\|\leq\epsilon\\ \|\boldsymbol{\delta}\|-\frac{\epsilon}{2}&\|\boldsymbol{\delta}\|>\epsilon\end{cases}~{}~{}.$
Since for any $\boldsymbol{\delta}$ it holds that $\|\boldsymbol{\delta}\|-\frac{\epsilon}{2}\leq g_{\epsilon}(\boldsymbol{\delta})\leq\|\boldsymbol{\delta}\|$, the solution of the relaxed problem Eq. (12) is at most $\frac{(n-1)\epsilon}{2}$ suboptimal with respect to that of the original problem defined in Eq. (11). Let us introduce a further auxiliary function,
$\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})=\begin{cases}\frac{\|\boldsymbol{\delta}\|^{2}}{2\epsilon}&\|\boldsymbol{\delta}^{\prime}\|\leq\epsilon\\ \frac{\|\boldsymbol{\delta}\|^{2}}{2\|\boldsymbol{\delta}^{\prime}\|}+\frac{\|\boldsymbol{\delta}^{\prime}\|-\epsilon}{2}&\|\boldsymbol{\delta}^{\prime}\|>\epsilon\end{cases}~{}~{}.$
The function $\hat{g}_{\epsilon}$ is a quadratic upper bound on $g_{\epsilon}(\boldsymbol{\delta})$ expanded at $\boldsymbol{\delta}^{\prime}$, thus $\forall\boldsymbol{\delta},\boldsymbol{\delta}^{\prime}:~{}\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})\geq g_{\epsilon}(\boldsymbol{\delta})$ and $\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta})=g_{\epsilon}(\boldsymbol{\delta})$. By examining each case we get the following,
$\hat{g}_{\epsilon}(\boldsymbol{\delta},\boldsymbol{\delta}^{\prime})$ $=$ $g_{\epsilon}(\boldsymbol{\delta}^{\prime})+\nabla{g_{\epsilon}}(\boldsymbol{\delta}^{\prime})^{\top}(\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime})+\frac{1}{2\,\max\big{(}\|\boldsymbol{\delta}^{\prime}\|,\epsilon\big{)}}\|\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime}\|^{2}$
$\leq$ $g_{\epsilon}(\boldsymbol{\delta})+\frac{1}{2\,\max\big{(}\|\boldsymbol{\delta}^{\prime}\|,\epsilon\big{)}}\|\boldsymbol{\delta}-\boldsymbol{\delta}^{\prime}\|^{2}~{}~{}.$
Hence, Algorithm 1 is a proximal method where we iteratively apply the multivariate subgradient-following of the previous section to obtain an exact solution for each adaptively constructed upper bound. We follow a similar skeletal proof to show the convergence.
We denote by $\eta^{t}=\min_{i}\rho_{i}^{t}\geq\epsilon$ and the optimum of the relaxed problem as $\hat{\phi}^{*}=\hat{\phi}({\mathbf{x}}_{1:n}^{*})$. We derive a series of bounds as the means to show convergence, starting with the definitions of $g_{\epsilon}$ and $\hat{g}_{\epsilon}$,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$ $=$ $\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}^{t+1}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}^{t+1}-{\mathbf{x}}_{i}^{t+1})$
$\leq$ $\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}^{t+1}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\hat{g}_{\epsilon}({\mathbf{x}}_{i+1}^{t+1}-{\mathbf{x}}_{i}^{t+1},{\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t})$
$=$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\hat{g}_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i},{\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t})$
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\frac{1}{2\rho_{i}^{t}}\|({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-({\mathbf{x}}_{i+1}^{t}-{\mathbf{x}}_{i}^{t})\|^{2}$
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\frac{1}{\rho_{i}^{t}}\big{(}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i+1}^{t}\|^{2}+\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\|^{2}\big{)}$
$\leq$ $\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})+\sum_{i=1}^{n}\frac{2}{\eta^{t}}\|{\mathbf{x}}_{i}-{\mathbf{x}}_{i}^{t}\|^{2}~{}~{}.$
Next we introduce a auxiliary sequence ${\mathbf{z}}_{1:n}(\mu)=(1-\mu)\,{\mathbf{x}}_{1:n}^{t}+\mu\,{\mathbf{x}}_{1:n}^{*}$ and upper bound the last inequality using a minimizer constrained to the line segment from ${\mathbf{x}}_{1:n}^{t}$ to ${\mathbf{x}}_{1:n}^{*}$. We then use the convexity of the objective. These two steps amount to,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$$\leq$ $\min_{\mu\in[0,1]}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{z}}_{i}(\mu)-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}g_{\epsilon}({\mathbf{z}}_{i+1}(\mu)-{\mathbf{z}}_{i}(\mu))+\sum_{i=1}^{n}\frac{2}{\eta^{t}}\|{\mathbf{z}}_{i}(\mu)-{\mathbf{x}}_{i}^{t}\|^{2}$
$\leq$ $\min_{\mu\in[0,1]}(1-\mu)\,\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\mu\,\hat{\phi}({\mathbf{x}}_{1:n}^{*})+\sum_{i=1}^{n}\Big{(}\frac{2\mu^{2}}{\eta^{t}}-\frac{\mu(1-\mu)}{2}\Big{)}\|{\mathbf{x}}_{i}^{t}-{\mathbf{x}}_{i}^{*}\|^{2}~{}~{}.$
Finally, we replace the minimizer w.r.t. $\mu$ with a specific choice $\mu=\frac{\eta^{t}}{4+\eta^{t}}$ and get,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})$$\leq$ $\frac{4}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{t})+\frac{\eta^{t}}{4+\eta^{t}}\hat{\phi}({\mathbf{x}}_{1:n}^{*})~{}~{}.$
Therefore, the optimality gap of the surrogate $\hat{\phi}$ satisfies,
$\hat{\phi}({\mathbf{x}}_{1:n}^{t+1})-\hat{\phi}^{*}\leq\frac{4}{4+\eta^{t}}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}\leq\frac{4}{4+\epsilon}\big{(}\hat{\phi}({\mathbf{x}}_{1:n}^{t})-\hat{\phi}^{*}\big{)}~{}~{}.$
In summary, we get that the sequence $\hat{\phi}({\mathbf{x}}_{1:n}^{t})$ converges linearly to $\hat{\phi}^{*}$.
To conclude the section we would like to underscore the computational advantage of using a surrogate loss with exact subgradient following over a dual ascent algorithm specifically tailored for the problem. The dual proximal gradient (DPG) is described and shortly analyzed in Appendix B. To do so, we generated synthetic data of varying levels of variational sparsity as follows. For each sparsity level, let $\mathcal{D}$ denote a distribution whose density function satisfies $\mathcal{D}({\mathbf{x}})\sim\text{e}^{-\lambda\|{\mathbf{x}}\|}$ such that its probability mass within the unit ball is equal to a prescribed fusion probability. We first sample a sequence of Boolean variables $\{\nu_{i}\}$ from the Bernoulli distribution with a predefined probability which designated the chance of fusing two consecutive vectors. When $\nu_{i}=1$, we sampled dual variables $\boldsymbol{\alpha}_{i}$ from $\mathcal{D}$ over the unit ball and set ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}=\mathbf{0}$. Otherwise, we sampled ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ from $\mathcal{D}$ outside the unit ball and set $\boldsymbol{\alpha}_{i}\in\partial\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|$. We then constructed $\mathbf{y}_{i}$ based on the optimality condition, $\mathbf{y}_{i}={\mathbf{x}}_{i}-\boldsymbol{\alpha}_{i}+\boldsymbol{\alpha}_{i-1}$. This randomized generation process ensures that ${\mathbf{x}}_{1:n},\boldsymbol{\alpha}_{0:n}$ are primal-dual optimal for $\mathbf{y}_{1:n}$.
We ran both DPG and the Untangled Lariat for 100 iterations each. We then used the resulting solution and merged two consecutive vectors when the $2$-norm of their difference was smaller than $10^{-8}$. For each level of variational sparsity we generated $100$ random sequences of vectors of dimension $d=100$ where each sequence is of length $n=100$. We measured the $2$-norm between the solution found by the algorithms and the primal-dual optimal solution which was used to generate the synthetic data. The results, shown in Figure 6, clearly indicate that Untangled Lariat is superior to DPG which despite the latter being tailored for the problem. The gap between ${\mathbf{x}}^{*}_{1:n}$ and $\hat{{\mathbf{x}}}_{1:n}$ found by Untangled Lariat is insensitive to the fusion probability whereas DPG’s performance deteriorates as the level of sparsity increases. Equally, if not more important, Untangled Lariat exhibits much superior recovery rate of fusion events and in fact seems to slightly improve with the level of sparsity. In contrast, DPG’s rate of recovery drastically falls to the point that at a sparsity level of $70\%$ almost no fusion events were identified.
Figure 6: Untangled Lariat vs. DPG for sparse multivariate discovery with a $2$-norm penalty.
5.2 Multivariate Regression With $\infty$-norm Variational Penalty
The second vector fusion penalty we consider is the infinity norm of the difference vectors ${\mathbf{x}}_{i}-{\mathbf{x}}_{i+1}$ which entails the problem of finding the optimal solution of,
$\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}~{}~{}.$ (13)
We associate a slack variable with each difference vector $\xi_{i}$ and rewrite the problem as,
$\min_{{\mathbf{x}}_{1:n},\boldsymbol{\xi}\geq 0}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}~{}.$
We cast the problem as a gradient-based search for the optimum $\xi_{1:n-1}^{*}$.
On each gradient step we solve a residual problem in ${\mathbf{x}}$ using the subgradient following procedure with bounded variation from the previous section. Formally, given any $\xi_{1:n-1}\geq 0$ we use the problem given by Eq. (13) to define $\mathcal{L}(\xi_{1:n-1})$ as follows,
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}$ s.t. $\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|_{\infty}\leq\xi_{i}~{}~{}.$
In order to compute the subgradient of $\mathcal{L}$, we introduce the dual variables and write the constrained problem as,
$\mathcal{L}(\xi_{1:n-1})=\min_{{\mathbf{x}}_{1:n}}\,\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}+\sum_{i=1}^{n-1}\boldsymbol{\alpha}_{i}^{\top}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-\xi_{i}\|\boldsymbol{\alpha}_{i}\|_{1}~{}~{}.$
From strong duality we can swap the order of min and max in the minimax problem to get,
$\mathcal{L}(\xi_{1:n-1})$ $=$ $\max_{\boldsymbol{\alpha}_{1:n-1}}\,\min_{{\mathbf{x}}_{1:n}}\;\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|^{2}+\sum_{i=1}^{n-1}\xi_{i}+\sum_{i=1}^{n-1}\boldsymbol{\alpha}_{i}^{\top}({\mathbf{x}}_{i+1}-{\mathbf{x}}_{i})-\xi_{i}\|\boldsymbol{\alpha}_{i}\|_{1}$
$=$ $\max_{\boldsymbol{\alpha}_{1:n-1}}\;\frac{1}{2}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\|_{2}^{2}+\sum_{i=1}^{n-1}\xi_{i}\big{(}1-\|\boldsymbol{\alpha}_{i}\|_{1}\big{)}~{}~{}.$
Denote by $\boldsymbol{\alpha}_{1:n-1}^{*}$ the maximizer of the above equation and let $\mathbf{u}$ be the vector whose $i$’th coordinate is $u_{i}=1-\|\boldsymbol{\alpha}_{i}^{*}\|_{1}$. For any $\xi_{1:n-1}^{\prime}\geq 0$, we have
$\mathcal{L}(\xi_{1:n-1})+\mathbf{u}^{\top}(\xi_{1:n-1}^{\prime}-\xi_{1:n-1})=$
$\displaystyle\frac{1}{2}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}^{*}-\boldsymbol{\alpha}_{i-1}^{*}\|_{2}^{2}+\sum_{i=1}^{n-1}\xi_{i}^{\prime}\big{(}1-\|\boldsymbol{\alpha}_{i}^{*}\|_{1}\big{)}\leq\mathcal{L}(\xi_{1:n-1}^{\prime})~{}.$
Therefore $\mathbf{u}\in\partial\mathcal{L}(\xi_{1:n})$, namely, it is a subgradient of $\mathcal{L}$ at $\xi_{1:n-1}$. Using strong duality again, we can obtain the optimal dual variables by first solving the primal problem and then infer the dual variables $\boldsymbol{\alpha}_{:}$ from optimality conditions. Given $\xi_{1:n-1}$ the primal is separable and we thus can apply the subgradient following from 3.5 to obtain both the optimal primal and dual variables. The slack variables $\xi_{1:n-1}$ are updated using gradient stepping. To do so, we derive upper bounds such that $\xi_{i}^{*}\leq D_{i}$ and then apply AdaGrad [9] with an initial per-coordinate learning rate of $D_{i}/\sqrt{2}$. The derivation of the upper bounds is given in Appendix C.
[…]
Figure 7: Untangled Lariat vs. DPG for sparse multivariate discovery using the the $\infty$-norm .
6 High-Order Variational Penalties
The differences $x_{i+1}-x_{i}$ can morally be viewed as discrete derivatives. The focus of this section is higher order discrete derivatives as variational penalties. For example, penalizing for the second order derivative amounts to,
$\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\lambda_{i}\big{|}(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})\big{|}~{}~{}.$
Rather than discussing specific forms of discrete difference operators, we analyze a general case stemming from convolution operators. Formally, let $\boldsymbol{\tau}=(-1,\boldsymbol{\bar{\tau}})$ where $\boldsymbol{\bar{\tau}}=(\tau_{1},\tau_{2},\ldots,\tau_{k})$ be vectors over $\mathbb{R}^{k+1}$ and $\mathbb{R}^{k}$ respectively. The convolution of ${\mathbf{x}}=(x_{1},\ldots,{\mathbf{x}}_{n})$ with $\boldsymbol{\tau}$ is a vector whose $i$’th entry is, $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=\sum_{m=0}^{k}\tau_{m}\,x_{i-m}$.
Equipped with these definitions we now focus on problems of the following general form,
$\min_{\mathbf{x}}\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=k}^{n-1}g_{i}\big{(}(\boldsymbol{\tau}\ast{\mathbf{x}})_{i+1}\big{)}~{}~{}.$ (14)
Since scaling is a degree of freedom we set $\tau_{0}=-1$ and use the definition above $\boldsymbol{\tau}=(-1,\boldsymbol{\bar{\tau}})$ to get that $(\boldsymbol{\tau}\ast{\mathbf{x}})_{i}=(\boldsymbol{\bar{\tau}}\ast{\mathbf{x}})_{i}-x_{i}$. Here we on purpose shift the indices of $g_{i}$ by $1$ to keep the same notation as used in the algorithms above when plugging in $\boldsymbol{\bar{\tau}}=(1)$. We use $\boldsymbol{\tau}\circledast{\mathbf{v}}$ to denote the convolution of a vector ${\mathbf{v}}$ with the reversal of $\boldsymbol{\tau}$ which we define as,
$(\boldsymbol{\tau}\circledast{\mathbf{v}})_{i}=\sum_{m=0}^{k}\tau_{m}v_{i+m}~{}~{}.$
We again confine ourselves to settings for which strong duality holds thus,
$U$ $=$ $\min_{{\mathbf{x}}_{1:k}}\,\max_{\alpha_{k}}\,\min_{x_{k+1}}\ldots\max_{\alpha_{n-1}}\,\min_{x_{n}}\>\sum_{i=1}^{n}h_{i}(x_{i})+\sum_{i=k}^{n-1}\left[\alpha_{i}(\boldsymbol{\tau}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})\right]$
Analogously, we define a sequence of intermediate functions,
$f_{i}({\mathbf{x}})$$\equiv$ $\max_{\alpha_{i}}\,\ldots\min_{x_{n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(\boldsymbol{\tau}*{\mathbf{x}})_{j+1}-g_{j}^{*}(\alpha_{j})\right]~{}~{}.$ (15)
This definition means that $f_{i}:\mathbb{R}^{i}\to\mathbb{R}$ and thus formally defined as $f_{i}({\mathbf{x}}_{1:i})$ since ${\mathbf{x}}_{i+1:n}$ is grounded to its optimal vector. As boundary conditions we set $f_{n}({\mathbf{x}})=0$. Similar to the constructions above we introduce an auxiliary function, $f_{i-\frac{1}{2}}({\mathbf{x}})=f_{i}({\mathbf{x}})+h_{i}(x_{i})$ which, like $f_{i}$, is a function of ${\mathbf{x}}_{1:i}$.
To simplify notation we denote by $\boldsymbol{\mathsf{x}}_{i}$ the prefix ${\mathbf{x}}_{1:i}$ and $\boldsymbol{\mathsf{a}}_{i}$ for $\boldsymbol{\alpha}_{i:n-1}$. Albeit the analogous notation, $\boldsymbol{\mathsf{x}}_{i}$ denotes a free variable whereas $\boldsymbol{\mathsf{a}}_{i}$ designates the optimal dual vector. Using these definitions we introduce a summation-free form for $f_{i}$,
$f_{i}({\mathbf{x}})$ $=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\bar{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+1}({\mathbf{x}})+h_{i+1}(x_{i+1})-\alpha_{i}x_{i+1}\right\}$ (16)
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\bar{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})+\min_{x_{i+1}}\left\{f_{i+\frac{1}{2}}(x_{i+1}|\boldsymbol{\mathsf{x}}_{i})-\alpha_{i}x_{i+1}\right\}$
$=$ $\max_{\alpha_{i}}\>\alpha_{i}(\boldsymbol{\bar{\tau}}*{\mathbf{x}})_{i+1}-g_{i}^{*}(\alpha_{i})-f_{i+\frac{1}{2}}^{*}(\alpha_{i}|\boldsymbol{\mathsf{x}}_{i})~{}~{},$
where $f_{i+\frac{1}{2}}(\cdot|\boldsymbol{\mathsf{x}}_{i})=f_{i+\frac{1}{2}}(\boldsymbol{\mathsf{x}}_{i},\cdot)$.
The subgradient conditions for locally-restricted optimality define two admissible sets,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial f_{i+\frac{1}{2}}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}~{}~{}.$
We next show that the following recursive form holds,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\bar{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(\mathsf{x}_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}$
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}~{}.$
Note that the admissible set consists of the full sequence of locally-admissible pairs. From strong duality, $S_{i}$ de-facto defines a point-to-set mapping from a prefix vector $\boldsymbol{\mathsf{x}}_{i}$ to its dual solutions of $f_{i}(\cdot)$. Hence, we can rewrite Eq. (15) as,
$f_{i}({\mathbf{x}})$ $=$ $\max_{\boldsymbol{\alpha}_{i:n-1}}\,\min_{{\mathbf{x}}_{i+1:n}}\>\sum_{j=i}^{n-1}\left[h_{j+1}(x_{j+1})+\alpha_{j}(\boldsymbol{\tau}*{\mathbf{x}})_{j+1}-g_{j}^{*}(\alpha_{j})\right]$
$=$ $\max_{\boldsymbol{\alpha}_{i:n-1}}\>\sum_{j=i}^{n-1}\left[\alpha_{j}(\boldsymbol{\tau}*\boldsymbol{\mathsf{x}}_{i})_{j+1}-h_{j+1}^{*}(-(\boldsymbol{\tau}\circledast\boldsymbol{\alpha})_{j})-g_{j}^{*}(\alpha_{j})\right]~{}~{}.$
Thus, given $\boldsymbol{\mathsf{x}}_{i}$, for any $\boldsymbol{\mathsf{a}}_{i}$ attaining the above maximum we have
$\forall m\in\{0,\ldots,k-1\}:\;(\boldsymbol{\bar{\tau}}\circledast\boldsymbol{\mathsf{a}}_{i})_{-m}\in\partial_{x_{i-m}}f_{i}(\boldsymbol{\mathsf{x}}_{i})~{}.$
Now, given $S_{i+1}$ we perform the following expansion,
$S_{i+\frac{1}{2}}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}\in\partial f_{i+1}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})+\partial h_{i+1}(x_{i+1})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i}):\mathsf{a}_{i}-(\boldsymbol{\bar{\tau}}\circledast\boldsymbol{\mathsf{a}})_{i}\in\partial h_{i+1}(x_{i+1})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i+1})\in S_{i+1}\big{\}}~{}~{}.$
Using the fact that $\mathsf{a}_{i}\in\partial f_{i+\frac{1}{2}}(\mathsf{x}_{i+1}|\boldsymbol{\mathsf{x}}_{i})~{}\Leftrightarrow~{}\mathsf{x}_{i+1}\in\partial f^{*}_{i+\frac{1}{2}}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})$ we get,
$S_{i}$ $=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}\in\partial f_{i+\frac{1}{2}}^{*}(\mathsf{a}_{i}|\boldsymbol{\mathsf{x}}_{i})+\partial g_{i}^{*}(\mathsf{a}_{i})\big{\}}$
$=$ $\big{\{}(\boldsymbol{\mathsf{x}}_{i},\boldsymbol{\mathsf{a}}_{i}):(\boldsymbol{\bar{\tau}}*\boldsymbol{\mathsf{x}})_{i+1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i})\;;\;(\boldsymbol{\mathsf{x}}_{i+1},\boldsymbol{\mathsf{a}}_{i})\in S_{i+\frac{1}{2}}\big{\}}~{}~{}.$
Finally, to satisfy the boundary conditions we need to find the solution of,
$\min_{{\mathbf{x}}_{1:k}}f_{k}({\mathbf{x}})+\sum_{i=1}^{k}h_{i}(x_{i})~{}~{}.$
The subgradient conditions imply that
[…]
where we used the above property to compute $\partial_{\mathsf{x}_{i}}f_{k}({\mathbf{x}})$.
Figure 8: Exact vs. approximate subgradient following for second-order variational penalties.
Second-order Variational Penalties.
To illustrate the power of the convolution-based representation, we provide a specific derivation for variational penalties which correspond to second order finite differences. To do so we simply need to set $\boldsymbol{\bar{\tau}}=(2,-1)$. Since $f_{i}(\cdot)$ depends only on $x_{i-1}$ and $x_{i}$, we can represent the subgradient sets as,
$S_{i}=\{\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})\;;\;\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})\}~{}\mbox{ and }~{}S_{i+\frac{1}{2}}=\{\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}~{}~{}.$
Assume this holds true for $S_{i+1}~{}\mbox{ i.e. }~{}S_{i+1}=\{\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})\;;\;\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})\}$. The definition of the recursive map for $i+\frac{1}{2}$ implies that,
$\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1})=2\mathsf{a}_{i+1}(\mathsf{x}_{i},\mathsf{x}_{i+1})-\mathsf{a}_{i+2}(\mathsf{x}_{i},\mathsf{x}_{i+1})+\mathsf{x}_{i+1}-y_{i+1}~{}\mbox{ and }~{}\mathsf{a}_{i+1}^{\prime}(\cdot)=\mathsf{a}_{i+1}(\cdot)~{}~{}.$
To obtain the recursive map $S_{i+\frac{1}{2}}\to S_{i}$, we need to extract $\mathsf{x}_{i+1}$ from the following subgradient set equation,
$2\mathsf{x}_{i}-\mathsf{x}_{i-1}-\mathsf{x}_{i+1}\in\partial g_{i}^{*}(\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},\mathsf{x}_{i+1}))~{}~{}.$
Similar to the fused lasso, the solution is given by the backward map $T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})$ defined as
$T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\Big{[}2\mathsf{x}_{i}-\mathsf{x}_{i-1}\Big{]}_{z_{i+1}^{-}(\mathsf{x}_{i})}^{z_{i+1}^{+}(\mathsf{x}_{i})}~{}~{},$
where $z_{i+1}^{\pm}(\mathsf{x}_{i})$ denotes the boundary values such that $\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},z_{i+1}^{\pm}(\mathsf{x}_{i}))=\pm 1$. Using this form of $T$ back in $\mathsf{x}_{i+1}$, we obtain
$\mathsf{a}_{i}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))$
and
$\mathsf{a}_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i})=\mathsf{a}_{i+1}^{\prime}(\mathsf{x}_{i},T_{i+1}(\mathsf{x}_{i-1},\mathsf{x}_{i}))~{}~{}.$
Here, $\mathsf{a}_{i}(\cdot)$ is a piecewise linear manifold in $\mathbb{R}^{2}$. We store triplets of the form $\{v_{j}^{1},v_{j}^{2},\alpha_{j}\}$ for vertices of each facet defining $\mathsf{a}_{i}$. We conjecture that the number of facets grows as $\mathcal{O}(n^{2})$ but leave formal analysis of runtime as future research. We empirically observed that the description of $\mathsf{a}_{i}$ is typically linear in the length of the input sequence. Alas, a disadvantage of the exact subgradient following described here is the need to maintain a data structure for high dimensional objects. We describe an alternative approximation yielding a simple iterative procedure for the problem in Appendix D.
We next demonstrate the effectiveness of exact subgradient following over the approximation-based method. To do so, we generated synthetic data of varying sequence length similar to the previous section. We sampled $(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})$ from $\mathcal{D}$ and then constructed $\mathbf{y}$ using the optimality conditions. Throughout the experiment, the sparsity level was set to $50\%$. In each experiment, we ran the approximation-based algorithm for $100$ iterations. The results are summarized in Figure 8. The advantage of exact subgradient following over approximation both in terms of runtime and optimality is apparent. The exact subgradient following weakly depends on the problem size whereas the quality of the approximate solution as well as run time quickly deteriorates even for short sequences.
7 Empirical Study
Prior to recapping the main results and concluding, we present in this section an empirical study that underscores the potential of the high-order Untangled Lariat. The experiments do not pit one version of algorithm versus another and claim superiority of a particular version, on the contrary. We describe experiments with three different financial tickers of different temporal characteristics in order to exhibit the merits of different versions in the light of different stochastic settings. The three datasets that we experimented with are SPX, VIX, and SFXRSA which we describe shortly below.
Figure 9: Comparison of Fused Lasso and Untangled Lariat on financial indices.
SPX is the ticker symbol for the Standard and Poor’s 500, a stock market index tracking the stock performance of 500 of the largest companies listed on stock exchanges in the United States. It is one of the most commonly followed equity indices and includes approximately 80% of the total market capitalization of US public companies. SPX is the case study of many research papers and has garnered numerous stock market models.
VIX is the ticker symbol for the Chicago Board Options Exchange’s CBOE Volatility Index, a popular measure of the stock market’s expectation of volatility based on S&P 500 index options. It is calculated and disseminated on a real-time basis by the CBOE, and is often referred to as the fear index or fear gauge. The VIX index tends to oscillate between periods of low volatility with a small magnitude of it and changes with spike-like transient periods. The VIX is rather difficult to model.
SFXRSA is a S&P Case-Shiller Home Price Index that measures the price level of existing single-family homes in San Francisco. The index reflects the average change in home prices in terms of percentage changes in housing market prices given a constant level of quality. Changes in the types and sizes of houses or changes in the physical characteristics of houses are specifically excluded from the calculations to avoid incorrectly affecting the index value. It is thus a slowly varying signal with some seasonal effect.
The three particular variants that we examined were the original Fused Lasso, and Untangled Lariat with a second order penalty equivalent for penalizing the total variation of the discrete second derivative. For comparison we also ran the Fused Lasso on the sequence of differences $y_{i+1}-y_{i}$. The latter mimics the $2$’nd order Untangled Lariat. The discrete derivative suppresses the low frequency components of the original sequence, obviating completely the zero frequency component. We tuned $\lambda$ for the three variants below so that their residual error of the entire series of each financial index is the same. Thus, the different behavior of each filtering approach reflects their characteristics and merits. As written above, the goal of the comparison is by no means to show superiority or a universal advantage of the nascent Untangled Lariat over the Fused Lasso or other methods mentioned in this paper. Indeed, there are settings in which the Fused Lasso would constitute the most viable data analysis tool, for instance during prolonged periods of economic stagnation.
The results for the three financial indices are depicted on the left column of Fig. 9. The three procedures seem to capture the overall tendency of the three different indices. The approximated indices generated by Fused Lasso are, as one would expect, piecewise constant. Overall, the approximation does not seem to provide an aesthetic refinement of the original data. We would henceforth more closely examine solely the performance of the Fused Lasso on the sequence of difference and the Untangled Lariat on the original data. Upon a first examination it is apparent that the Fused Lasso on the difference sequences is capable of capturing transient periods during which there is typically a sharp upturn followed by a similar or less pronounced downturn. However, there also seems to be periods of systematic bias within the approximated sequences. Since the finite derivatives of the original indices eliminates zero frequency effects, the biased estimates do not affect the overall error.
To further qualitatively examine the results, we provide on the right column of Figure 9 zoomed-in snippets of an upturn-downturn period. These snippets are provided for visualization purposes and were not handpicked. All three variants were provided the entire sequence, not the particular snippets we visualize. Traversing the figures from top hand-side to bottom hand-side, for SPX we see that the Fused Lasso’s approximation is over-fragmented and does not provide a faithful approximation of the up-then-down characteristics. The Fused Lasso on the sequence of price changes does capture nicely the temporal wedge shape but it oversubscribes to small changes on occasion. The $2$’nd order variant of Untangled Lariat nicely captures the wedge characteristics with about $4$ piecewise linear segments.
By construction and definition of the VIX index, it is more volatile than SPX. In the middle row we examine a snippet of the VIX index during a morally up-down period albeit with further nuanced changes that could be significant for a financial data analyst. The approximation of the Fused Lasso on the price changes in comparison with $2$’nd order Untangled Lariat underscores a natural trade-off. The latter approximates VIX by morally 4 linear segments, a period without a substantial change that is approximated as a sequence of fixed values, followed by a single segment of upturn, concluding with 3 segments of linear decrease in the index. It thus faithfully captures the overall trend. In contrast, the Fused Lasso on price changes still captures the nature of the sub-sequence is more fragmented and sensitive to local changes albeit resulting with the same approximation error as the Untangled Lariat due to the respective choices of $\lambda$. It thus provides a more pinpointed tool during high volatility periods that are absent from the overall market trend. Lastly, for the real-estate index SFXRSA, all three variants seem to capture quite faithfully the overall behavior of the slowly changing real-estate market in San Francisco.
We would like to note that we can possibly gain the best of all approaches with a further generalized version of Untangled Lariat with composite variational penalty consisting of two terms such as,
$\lambda_{1}|x_{i+1}-x_{i}|+\lambda_{2}|x_{i-1}-2x_{i}+x_{i+1}|~{}~{}.$
We leave the derivation and analysis of Untangled Lariat with composite variational penalties to future research.
8 Concluding Remarks
This work (subjectively) makes the following novel contributions:
  • Provides a unified approach for sequence approximation w.r.t. Bregman divergences with general total variation penalties.
  • Describes and analyzes a novel subgradient following procedure, Untangled Lariat, with succinct code that can be readily used with existing and new variation penalties.
  • Untangled Lariat entertains the same time complexity as known algorithms, each of which designed (by different authors) for a concrete variational penalty.
  • Introduces and analyzes a multivariate generalization where each element of the sequence is a vector. The base case of squared $2$-norm as the variational penalty of two consecutive vectors bears a closed form for subgradient following.
  • Derives iterative algorithms for sparse multivariate total variation using the $2$-norm or the $\infty$-norm by employing the norm-squared version as surrogate.
  • Describes and analyzes nascent high-order Untangled Lariat with poly-time subgradient following.
There are several possible extensions of the framework presented in this paper. The univariate penalties are readily extendable to non-symmetric counterparts and the addition of insensitivity regions. For instance, a generalization the symmetric penalty of the fused lasso amounts to,
$\lambda_{i}^{\mathsf{R}}\big{[}x_{i+1}-x_{i}-\epsilon_{i}^{\mathsf{R}}\big{]}_{+}\,+\,\lambda_{i}^{\mathsf{L}}\big{[}x_{i}-x_{i+1}-\epsilon_{i}^{\mathsf{L}}\big{]}_{+}~{}~{}.$
By choosing $\lambda_{i}^{\mathsf{R}}=\lambda_{i}^{\mathsf{L}}=\lambda_{i}$ and $\epsilon_{i}^{\mathsf{R}}=\epsilon_{i}^{\mathsf{L}}=0$ we obtain the fused lasso’s penalty. To encompass this version we simply need to compute $\partial g_{i}^{*}$ from which we derive a new mapping $\mathsf{x}\mapsto\mathsf{a}_{i}(\mathsf{x})$. The subgradient procedure uses a similar form for the mapping given in Eq. (7) though the introduction of insensitivity regions to the left and/or the right of the loci where $x_{i}=x_{i+1}$ would introduce two additional inflection points. The resulting procedure for constructing the subgradient would require at most $\mathcal{O}(n\log n)$ time.
We limited our discussion to sequential penalties which can described as a chain of local connectivities $x_{1}\leftrightarrow x_{2}\leftrightarrow\ldots x_{i}\leftrightarrow x_{i+1}\ldots$ forming a left-to-right dependency graph. Kolmogorov et. al [18] examined a setting where the total variation penalties form a tree. Denote the variation dependency graph by $G=(V,E)$ where $V=[n]$ and $(i,j)\in E$ if there exists a penalty term for the variables $x_{i}$ and $x_{j}$. Using the graph representation, the penalized problem takes the form of,
$U=\min_{{\mathbf{x}}}\>\sum_{i}\frac{1}{2}(x_{i}-y_{i})^{2}+\sum_{(i,j)\in E}|x_{i}-x_{j}|~{}~{}.$
In settings for which strong duality holds the above problem can be re-parameterized as,
$U$ $=$ $\max_{\|\boldsymbol{\alpha}\|_{\infty}\leq 1}\,\min_{{\mathbf{x}}}\>\frac{1}{2}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}+\sum_{(i,j)\in E}\alpha_{i,j}(x_{i}-x_{j})~{}~{}.$
When $E$ defines a tree the subgradient following procedure can utilize the form described in [18] albeit increasing the time complexity to $\mathcal{O}(n^{2})$.
Acknowledgments
We would like to thank Tomer Koren and Yishay Mansour for fruitful discussions. The short discussion above is motivated by conversations with the late Prof. Tali Tishby who is missed dearly. This manuscript was written without the usage of AIML tools.
Appendix A Improved Barrier Lariat using Skiplists
We store each $\mathsf{a}_{i}(\cdot)$ as a linked-list data structure of implicit change points (vertices) and explicit linear intervals connecting two consecutive vertices. A vertex $(v_{j},\mathsf{a}_{i}(v_{j}))=(v_{j},\alpha_{j})$ represents a point where $\mathsf{a}_{i}(\cdot)$ changes slope. An edge $(d_{j},s_{j})$ stores the length and slope of the line segment between $v_{j}$ and $v_{j+1}$, thus we have,
$d_{j}=v_{j+1}-v_{j}\,;\,s_{j}=(\alpha_{j+1}-\alpha_{j})/d_{j}~{}~{}.$
This representation allows lazy evaluations in a similar fashion to the fused lasso. The Kvetsh step does not change the lengths of intervals except for the new segment encapsulating the zero crossing. Since the length of all other line segments remain intact this edge-based representation facilitates an efficient implementation.
By construction of $\mathsf{a}_{i+1}$ we know that $\exists j$ for which $v_{j}=z_{i+1}-\lambda_{i+1}$ and $\mathsf{a}_{i+1}(z_{i+1}-\lambda_{i+1})=0$. This implies that in an explicit representation of $\mathsf{a}^{\prime}_{i}(\cdot)$ we would need to update vertex $j$,
$(z_{i+1}-\lambda_{i+1},0)~{}~{}\mapsto~{}~{}(z_{i+1}-\lambda_{i+1},z_{i+1}-\lambda_{i+1}-y_{i+1})~{}~{}.$
In the explicit view of the lazy representation of the ensuing edge, the slope $s_{j}$ of the segment $[v_{j},v_{j+1}]$ becomes $s_{j}+n-i$. We next find the zero crossing point of $\mathsf{a}^{\prime}_{i}(\cdot)$ by traversing the list of edges towards zero and locating the index $k$ for which $\alpha_{k}\leq 0$ and $\alpha_{k+1}>0$. In case such an index does not exist the zero of $\mathsf{a}^{\prime}_{i}(\cdot)$ is right of its rightmost vertex. Once we identify the enclosing interval, we calculate $z_{i}$ by solving a linear equation from $(v_{k},\alpha_{k})$ and $(v_{k+1},\alpha_{k+1})$. We next Kvetsh the list by inserting two vertices, $(z_{i}-\lambda_{i},0)$ and $(z_{i}+\lambda_{i},0)$, between $k$ and $k+1$, and update $d_{k}$ accordingly. This sequential traversal of the list in the worst-case would require linear time and thus total run time would remain intact at $\mathcal{O}(n^{2})$.
In order to employ a more efficient procedure for finding zero crossings, we make use of a skip-list data structure [20]. A skip-list is an extension of a linked list with multiple layers and with which the expected search time is $\mathcal{O}(\log{n})$. The skip-list has ideally in our setting $\mathcal{O}(\log{n})$ layers where layer $k$ contains all vertices whose indices modulo $2^{k}$ are zero. The search procedure starts from the top layer and proceeds downwards once we reach the interval encapsulating zero. In practice, a skip-list is a randomized data structure where each vertex appears in a higher layer with probability of $\frac{1}{2}$ to support efficient insertions.
We follow the same lazy representation as above. We store $(d_{j}^{k},s_{j}^{k})=(v_{j+1}^{k}-v_{j}^{k},s_{j}^{k})$ for edge $j$ at layer $k$. We also add a precursor vertex at the beginning of each layer. This vertex contains the smallest possible value of $\{v_{j}\}$ and serves as a boundary point. To construct $\mathsf{a}^{\prime}_{i}(\cdot)$ we update $\alpha_{0}\leftarrow\alpha_{0}+v_{0}-y_{i+1}$. Then, we follow the search procedure above until reaching the bottom layer. We calculate $z_{i}$ at the bottom layer by solving the same linear equation as above. During the insertion of vertex $(z_{i}-\lambda_{i},0)$, we flip a coin to decide whether it should appear in a higher layer. Note that along the searching path, we have calculated $(v_{j}^{k},\alpha_{j}^{k})$ and $(v_{j+1}^{k},\alpha_{j+1}^{k})$ per layer which contains the locus of the zero crossing. Thus, whenever a vertex should appear in a prior (higher) layer, we use the corresponding values for updating the connecting edges. We perform an equivalent update for $(z_{i}+\lambda_{i},0)$. Last, we update the $v_{0}\leftarrow v_{0}-\lambda_{i}$. The end result is an $\mathcal{O}(n\log{n})$ amortized time algorithm.
Appendix B Dual Proximal Gradient
We give here further details on the dual proximal gradient method (DPG) used in the empirical studies. DPG merits further investigation on its own and to the best of our knowledge the short analysis with general norm penalties below was not derived in previous research.
[…]
$\min_{{\mathbf{x}}_{1:n}}~{}\frac{1}{2}\sum_{i=1}^{n}\big{\|}{\mathbf{x}}_{i}-\mathbf{y}_{i}\big{\|}_{2}^{2}\;+\;\sum_{i=1}^{n-1}\big{\|}{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\big{\|}~{}~{}.$
The dual of this problem is,
$\max_{\boldsymbol{\alpha}_{0:n}}~{}\sum_{i=1}^{n}\big{\|}\mathbf{y}_{i}\big{\|}_{2}^{2}-\big{\|}\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\big{\|}_{2}^{2}~{}\;\text{s.t.}\;~{}\big{\|}\boldsymbol{\alpha}_{i}\big{\|}_{*}\leq 1~{}~{}.$
For boundary conditions we set $\boldsymbol{\alpha}_{0}\equiv\boldsymbol{\alpha}_{n}\equiv\mathbf{0}$. Denote by $B_{*}=\{{\mathbf{x}}:\|{\mathbf{x}}\|_{*}\leq 1\}$ the unit ball w.r.t the dual norm. Let $\Pi_{B_{*}}(\cdot)$ denote the projection operator onto the unit ball. In each iteration of dual proximal gradient,
$\boldsymbol{\alpha}_{i}\,\leftarrow\,\Pi_{B_{*}}\Big{(}\boldsymbol{\alpha}_{i}-\eta(\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}+\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i+1}-\mathbf{y}_{i+1})\Big{)}~{}~{}.$
To underscore the advantage of the sub-gradient following method, let us briefly examine the condition number of the dual problem. Denote by $\mathsf{D}$ the difference matrix
$\mathsf{D}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\end{bmatrix}\in\mathsf{R}^{n-1\,\times\,n}~{}~{}.$
The condition number of the dual problem is the same as that of $\mathsf{D}\mathsf{D}^{\top}$. If we choose $\boldsymbol{\alpha}$ such that
$\alpha_{i}=\frac{{(-1)}^{i}}{\sqrt{n-1}}~{}~{},$
we have $\|\mathsf{D}^{\top}\boldsymbol{\alpha}\|_{2}^{2}=\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=4-\frac{2}{n-1}$. Assume without loss of generality that $n$ is even and let us choose next
$\alpha_{i}=\frac{\frac{n}{2}-|i-\frac{n}{2}|}{\sqrt{n^{3}/12}}~{}~{},$
for which we have $\|\boldsymbol{\alpha}\|_{2}\geq 1$ and $\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=12/n^{2}$. Therefore, for $n\geq 3$ the largest eigenvalue of $\mathsf{D}\mathsf{D}^{\top}$ is at least $3$ whereas the smallest eigenvalue is at most $\mathcal{O}(n^{-2})$. Therefore, the condition number of the dual problem is $\Omega(n^{2})$ which renders dual ascent algorithms, and in particular DPG, slow to converge.
Appendix C Upper Bounds on Slack Variables
Denote the optimum of the original problem by ${\mathbf{x}}_{1:n}^{*}$, we claim that $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\leq\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$ and thus we can set $D_{i}=\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$. If $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}=0$, the bound clearly holds. Otherwise let us define,
$S=\{j:|x_{i+1,j}^{*}-x_{i,j}^{*}|=\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\}~{}~{}.$
Consider $j\in S$ and let $s_{i,j}=\text{sign}(x_{i+1,j}^{*}-x_{i,j}^{*})$. From the optimality condition, there exists $\boldsymbol{\alpha}_{1:n-1}$ such that
$x_{i+1,j}^{*}-x_{i,j}^{*}$ $=$ $(y_{i+1,j}+\alpha_{i+1,j}-\alpha_{i,j})-(y_{i,j}+\alpha_{i,j}-\alpha_{i-1,j})$
$=$ $y_{i+1,j}-y_{i,j}+\alpha_{i+1,j}-2\alpha_{i,j}+\alpha_{i-1,j}~{}~{}.$
In addition, we have $s_{i,j}\alpha_{i,j}\geq 0$. Multiplying both side by $s_{i,j}$ we get
$|x_{i+1,j}^{*}-x_{i,j}^{*}|$ $=$ $s_{i,j}(y_{i+1,j}-y_{i,j})+s_{i,j}\alpha_{i+1,j}+s_{i,j}\alpha_{i-1,j}-2|\alpha_{i,j}|$
$\leq$ $|y_{i+1,j}-y_{i,j}|+|\alpha_{i+1,j}|+|\alpha_{i-1,j}|-2|\alpha_{i,j}|~{}~{}.$
Summing over $j\in S$ and dividing by $|S|$ we get,
$\big{\|}{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\big{\|}_{\infty}$$\leq$ $\frac{1}{|S|}\Big{(}\sum_{j\in S}|y_{i+1,j}-y_{i,j}|+\|\boldsymbol{\alpha}_{i+1}\|_{1}+\|\boldsymbol{\alpha}_{i-1}\|_{1}-2\Big{)}$
$\leq$ $\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}~{}~{},$
where for obtaining the second inequality we used the fact that for $j\notin S:\alpha_{i,j}=0$ and $\|\boldsymbol{\alpha}_{i}\|_{1}=1$ since $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\neq 0$.
Appendix D Iterative Algorithm for High-Order Variational Penalties
Note that the following modified smooth problem can be solved in linear time,
$\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\lambda_{i}\big{|}(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})\big{|}^{2}~{}~{}.$
Since the problem is differentiable everywhere, the optimality condition amounts to a linear system $\mathsf{D}{\mathbf{x}}=\mathbf{y}$ where $\mathsf{D}$ is a five-diagonal matrix. The system can be solved using Gauss elimination to reduce the matrix to an upper triangular one in linear time. Once triangulated, ${\mathbf{x}}$ is simply read out from the resulting system. Then, similar to Algorithm 1, we iteratively replace the non-smooth objective by its quadratic upper bound. Namely, on iteration $t$, we obtain an exact solution from,
${\mathbf{x}}^{t+1}=\displaystyle\operatorname*{\arg\!\min}_{\mathbf{x}}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}{\big{|}(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})\big{|}^{2}}\,/\,{\rho_{i}^{t}}~{}~{},$
where $\rho_{i}^{t}=\displaystyle\max\big{(}\big{|}(x_{i+2}^{t}-x_{i+1}^{t})-(x_{i+1}^{t}-x_{i}^{t})\big{|},\,\epsilon\big{)}$.
The line of proof as above can be carried out except that now we have,
$|(x_{i+2}-2x_{i+1}+x_{i})-(x^{\prime}_{i+2}-2x^{\prime}_{i+1}+x^{\prime}_{i})|^{2}\leq 4\left(|x_{i+2}-x^{\prime}_{i+2}|^{2}+2|x_{i+1}-x^{\prime}_{i+1}|^{2}+|x_{i}-x^{\prime}_{i}|^{2}\right)~{}~{}.$
Summing over $i$ we end up with a factor of 16 instead of 4. This procedure can be generalized to $k$’th order finite difference penalties. Concretely, let us define the matrix by $\tilde{\mathsf{D}}$ the discrete differentiation matrix with zero padding,
$\tilde{\mathsf{D}}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\\ 0&0&0&\dots&0&0\\ \end{bmatrix}\in\mathsf{R}^{n\,\times\,n}~{}~{}.$
We also employ this matrix sans the last row in Appendix B. The penalized $k$’th order finite differences we care to solve is,
$\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\big{\|}{\mathbf{x}}-\mathbf{y}\big{\|}^{2}+\big{\|}\big{[}\mathsf{\tilde{D}}^{k}{\mathbf{x}}\big{]}_{1:n-k}\big{\|}_{1}~{}~{}.$
In terms of run time, we “pay” a factor of $4^{k}$ for the number of iterations required to converge and each iteration takes $\mathcal{O}(k^{2}n)$ for solving the linear system. In comparison, when using the dual proximal gradient method, the iteration complexity grows as $\mathcal{O}(\gamma^{k})$ where $\gamma=\Omega(n^{2})$ is the condition number of the first order problem. This implies a much inferior time complexity of $\mathcal{O}(n^{2k})$ for DPG. Pseudocode for $k>0$ is provided below.
Algorithm 2 $k^{\text{th}}$-Order Untangled Lariat
  initialize ${\mathbf{x}}^{0}=\mathbf{y}$
  for $t=1\ \textbf{to}\ T$ do
     $\mathsf{H}^{t}=\begin{bmatrix}\operatorname{diag}\bigg{(}\bigg{[}\Big{|}\big{[}\mathsf{\tilde{D}}^{k}\,{\mathbf{x}}^{t}\big{]}_{1:n-k}\Big{|}\bigg{]}_{\epsilon}\bigg{)}^{-1/2}&\mathsf{O}_{n-k\times k}\\ \mathsf{O}_{k\times n-k}&\mathsf{O}_{k\times k}\\ \end{bmatrix}$
     ${\mathbf{x}}^{t+1}=\displaystyle\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\frac{1}{2}\|\mathsf{H}^{t}\,\mathsf{\tilde{D}}^{k}\,{\mathbf{x}}\|^{2}$
  end for
  return $\mathbf{x}$
Figure 10: Python code of algorithms from Sec. 2.